### v1

# Robust Weight Imprinting: Insights from Neural Collapse and Proxy-Based Aggregation

Abstract
The capacity of a foundation model allows for adaptation to new downstream tasks. Weight imprinting is a universal and efficient method to fulfill this purpose. It has been reinvented several times, but it has not been systematically studied. In this paper, we propose a framework for imprinting, identifying three main components: generation, normalization, and aggregation. This allows us to conduct an in-depth analysis of imprinting and a comparison of the existing work. We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization. We determine those proxies through clustering and propose a novel variant of imprinting that outperforms previous work. We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time. Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes.
1 Introduction
In machine learning applications, training models from scratch is often not viable due to limitations in data and compute. A popular solution is to apply transfer learning [1, 2] based on foundation models (FMs) [3] that are pre-trained on a large amount of data. A common approach in practice to adapt an FM to a novel task is to freeze its parameters and replace the output layer with a new head, e.g., for classification.
Imprinting.
Qi et al. [4] propose a simple solution for few-shot classification, called imprinting. Namely, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean. These class means are representatives of the classes, which we generally call proxies. This results in an efficient method without the need for gradient-based optimization. A plethora of studies have emerged surveying this technique by adding complexity and adaptability [5, 6, 7, 8, 9, 10, 11]. Despite many adaptations, imprinting lacks a systematic comparison that unifies them. Understanding its variations could unlock greater efficiency and performance across many fields, making the method even more versatile and impactful.
Framework.
We present a unifying framework that enables a systematic comparison of existing imprinting techniques. More precisely, we can generalize prior work by decomposing imprinting into three principal steps (see fig. 1). During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (multiple proxies). Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors. Aggregation (AGG) entails the computation of the final output, e.g., a class label.
The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments. We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization.
Neural Collapse.
We investigate a recently discovered phenomenon called neural collapse [12], which provides a compelling explanation for why imprinting works. According to this phenomenon, when neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means [12, 13]. Our investigation proves that a measurement of neural collapse provides insights about imprinting.
Contributions.
In summary, our main results and contributions are:
  • We deconstruct weight imprinting into a framework composed of generation, normalization, and aggregation, and discuss variations for each of them, identifying prior work as special cases (section 3). To the best of our knowledge, we are the first to conduct a comprehensive analysis of imprinting to this scale (section 5).
  • We present a new imprinting method utilizing $k$-means clustering for weight generation (section 6.1) and show its benefits in certain few-shot scenarios (section 6.2).
  • To the best of our knowledge, we are the first to identify a connection between imprinting success and measures of neural collapse (section 6.3).
We make the source code to reproduce our results publicly available. [^1] https://github.com/DATEXIS/multi-imprinting/
Figure 1: Overview of our imprinting framework. The foundation model FM is frozen and shows neural collapse. The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$. The final output of the test data in T is computed by an aggregation (AGG) mechanism. Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively. During inference, embeddings are normalized according to NORMinf (not shown here).
2 Related Work
Imprinting and Few-Shot Learning.
Weight imprinting was introduced in [4] for the few-shot learning scenario. It is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples. Qi et al. [4] find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms. However, we are not limiting the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency. Imprinting has also been applied to object detection [5, 6], multi-label classification [8], semantic segmentation [10], and in combination with an attention mechanism to generate weights for the novel classes in a few-shot classification task [11].
Hosoda et al. [14] apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights. We consider this as one normalization scheme in our framework. Zhang et al. [15] apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available. They speculate whether normalization is a constraint in their imprinting model.
Before the era of deep learning, Mensink et al. [16] analyze the transferability of hand-crafted image features. They use a "nearest class multiple centroids" (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm. In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm. Our work, on the other hand, highlights efficient transfer learning provided by foundation models.
Transfer Learning.
Our work is related to the use of embedding vectors extracted with pre-trained models, which is one of the most straightforward transfer learning techniques, since the seminal works in computer vision [17] and natural language processing [18]. Kornblith et al. [19] showed that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks. In addition, Huh et al. [20] provided insights into the required quality of pre-training data. Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
Continual Learning (CL).
Class means have also been used as proxies in CL. Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL. Rebuffi et al. [21] dynamically select a subset of examples for each class and update internal representations via gradient descent. They use a nearest mean classifier (NMC) with respect to the saved examples. Janson et al. [22] use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings. However, they do not investigate the effect of normalization and using multiple proxies.
Findings of [23] show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored. In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
Neural Collapse (NC).
The phenomenon of NC was identified by [12] and refers to the convergence of the last-layer weight vectors to class means. It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) [13, 24, 25], but complete neural collapse is practically unrealistic [26]. In transfer learning, Galanti et al. [27] show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios. In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
3 Imprinting Framework
In order to find out how to best set the classifier weights of a foundational model in downstream tasks T, we create a framework (see fig. 1) that encompasses many different combinations, all of which work without gradient-based training. Thereby, we can unify all the existing imprinting strategies described in section 2.
We analyze multi-class classification scenarios in that we do not separate into base and new classes, but focus on all classes in a novel T at the same time. To investigate the effect of the number of samples given, we look at n-shot ($n\in\mathbb{N}$) scenarios. For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the regime of few-shot learning.
Overview.
We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG). The framework depicted in fig. 1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted. The FM remains frozen throughout the experiments. It receives data from T as inputs and produces embedding vectors. The training process generates weight vectors for each of the $c$ classes in T. Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre. The generated weight vectors per class are called proxies, prototypes, or representatives [28, 29, 30, 31]. These proxies are normalized according to NORMpost. As in [4], we do not use bias values. To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
Special Imprinting Cases.
Previously proposed imprinting methods can be defined as a special case of our framework. Figure 4 in section 6.1 provides an overview of existing imprinting strategies listed in literature for foundational models and benchmarks them with the new best-performing one we find through our framework. In total, we inspect all possible combinations (including variations in models, tasks, and seeds).
Weight Generation (GEN).
The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights. In contrast to [4] which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in [16]. This enables non-linear classification. We denote the number of proxies as $k$, ranging between $1$ and the number of samples. We investigate the following operations conducted per class to generate its proxies:
  • all: All embeddings (denoted as $k=\text{all}$).
  • k-random: $k$ random embeddings.
  • mean: The mean of all embeddings.
  • k-means: $k$-means cluster centers. $k=1$ is the same as mean.
  • k-medoids: $k$-medoids cluster centers.
  • k-cov-max (covariance-maximization): Top $k$ embeddings by covariance.
  • k-fps (farthest-point sampling): Iteratively selecting $k$ embeddings, such that it maximizes the distance from already selected ones (starting with random sample).
We choose this diverse list of methods to cover a wide range of approaches, ranging from heuristics (e.g., k-fps) to more complex algorithms (e.g., k-means). Note that only mean and k-means generate proxies that are not included in the given samples.
Normalization (NORM).
The main reason for applying normalization is to allow each embedding and weight vector to contribute equally on the same scale. The modes we allow are no normalization (none), $L^{2}$ normalization (L2), and quantile normalization (quantile).
L2 normalization can be applied to embeddings before GEN via NORMpre, to the generated weights via NORMpost, and to embeddings in inference via NORMinf. In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean $\|\cdot\|_{2}$ length.
quantile normalization [32, 33] can only be applied to generated weights. This non-linear operation distributes weights equally. Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class, and the reference distribution changes accordingly. In particular, for the first class there is no reference distribution to map to. This is different from [14], where new weights are matched to the distribution of the original classifier weights of the FM. Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
Aggregation (AGG).
There are various ways to use the generated weights per class during inference, especially when $k>1$. We focus on two different modes, max and m-nn. The former, max, computes the inner product of the input embedding and the imprinted weights and outputs the class label with the maximum activation. The latter, m-nn, uses the class weights as keys and the embeddings as values, and chooses the final winning output class via the $m$-nearest neighbor algorithm. The m-nn voting is weighted by the inverse of the distances to their nearest neighbor, turning it into weighted majority voting.
Note that max is the same as 1-nn in the case of L2 for NORMpost, since for any fixed embedding vector $v$ and variable proxy $w$, the argmin of $\|v-w\|^{2}=\|v\|^{2}-2\langle v,w\rangle+\|w\|^{2}$, calculated by 1-nn, is the same as the argmax of the inner product $\langle v,w\rangle$ calculated in max.
4 Measurement of Neural Collapse
Neural collapse (NC) [12] refers to the phenomenon that occurs on the last-layer classifier weights of neural networks in the terminal phase of training (TPT). When the network is trained well beyond zero training error, the learned embeddings of each class, assuming balanced classes, collapse to their class means. These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability. This results in an optimal linearly separable state for classification. In fig. 2 (left), we illustrate the collapse of a FM on its pre-training data. The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).
Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification. We focus on variability collapse ($\mathcal{NC}_{1}$), using the equation 1 from [13]:
$\mathcal{NC}_{1}=\tfrac{1}{c}\mathrm{trace}(\Sigma_{W}{\Sigma^{+}_{B}}),$ (1)
where $\Sigma_{W},\Sigma_{B}\in\mathbb{R}^{l\times l}$ are within- and between-class covariance matrices, respectively, $l$ is the dimension of the embedding vector, $c$ is the number of classes, and + symbolizes the pseudo-inverse. Based on the equation, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse. In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig. 11). Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $c$. According to NC, imprinting the mean, as originally done in [4], is best when $\mathcal{NC}_{1}$ is small. We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class. We investigate this in section 6.3.
Figure 2: Left: The embeddings of the pre-training data, after being used to train the foundation model FM, show neural collapse, as each class ($o_{1},\dotsc,o_{4}$) is evenly separated in space and accumulates around their respective class means. Right: For a novel task with classes $c_{1},c_{2}$ (pink and brown) scatter around the collapsed pre-trained classes (gray).
5 Experimental Setup
Models.
We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture. In neural collapse investigations (section 6.3), we also use resnet50 [34] and swin_b [36]. All of them are pre-trained on ImageNet-1K (ILSVRC 2012) [37]. To generate the embeddings, we use PyTorch’s torchvision models.
Tasks.
To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10  [^2] https://www.cs.toronto.edu/~kriz/cifar.html, each containing 10 classes. We mainly focus on the three T containing all ten classes. Furthermore, we look at smaller tasks only containing classes $\{0,1,2\}$, and the two tasks containing classes $\{1,3,5,7,9\}$ resp. $\{0,2,4,6,8\}$. This random selection of $3\cdot 4=12$ tasks adds variation to our evaluation.
In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet). As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations. For ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis. These tasks are called "$d$ in $1$", $d=1,\dotsc,10$, each containing $10$ different labels. More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels. See fig. 3 for a simplified example. We do this random sampling 10 times which results in $10\cdot 10=100$ tasks.
Figure 3: Combining multiple classes into one to create tasks with multi-modal class distributions. Example: "$d$ in 1", $d=2,3$.
Scale.
In total, we run approximately 350 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds. This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
Evaluation.
Throughout our experiments, the median accuracy on the test set for three different seeds is reported, if not otherwise specified. In sections 6.1 and 6.2, we investigate the imprinting performance by varying the FM (2) and T (12). We then sort the combinations by their final accuracy. There are $2\cdot 12=24$ potentially different ranks for each of the combinations. We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [40]. [^3] Code used to generate these diagrams can be found here [41]. In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy. We consider differences significant if $p<0.05$.
In experiments with neural collapse (section 6.3), we investigate four FMs on 100 ImageNet tasks and the three tasks containing all of MNIST, FashionMNIST, and CIFAR-10, respectively.
6 Results
Our main experimental insights are:
  1. 1.
    Our imprinting framework generalizes previous methods, and we find a new superior imprinting strategy (section 6.1).
  2. 2.
    We show that our strategy is beneficial in few-shot scenarios with as little as 50 samples per class (section 6.2).
  3. 3.
    We identify a correlation between imprinting success utilizing multiple proxies and measures of neural collapse (section 6.3).
6.1 Best Imprinting Strategy
Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
Qi et al. [4] L2 mean L2 L2 max 87.72
Hosoda et al. [14] none mean quantile none max 82.15
Janson et al. [22] none mean none none 1-nn 87.65
Ours L2 k-means L2 L2 max 91.47
[…]
[table]A table beside a figure
Figure 4: Previously studied imprinting strategies are special cases within our framework. The framework enables the creation of a novel configuration ("Ours") that outperforms previous work across FMs and Ts by a large margin with statistical significance. Here, $k=20$ is chosen.
We provide a comparison between memory-constrained methods used for imprinting on foundation models in fig. 4, namely, [4, 14, 22], as well as a novel configuration that results from our framework which we call “Ours”. We investigate the impact of using m-nn aggregation on all data afterward. We focus on $k=20$ and find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all other approaches by a margin of $4\%$ with statistical significance, as can be seen in the CD diagram. Below, we analyze each of the components of the framework separately.
Weight Generation (GEN).
To analyze the impact of GEN, we first focus on the max aggregation. For the weight generation analysis, we do not fix NORM, but simply show the run with the best NORM combination, if not otherwise specified. The m-nn aggregation and different values for NORM are analyzed later in this section.
GEN $k$ Avg. acc. %
k-means 20 91.95
k-medoids 20 87.31
mean 1 87.76
k-cov-max 20 84.74
k-random 20 82.79
k-fps 2 63.90
[…]
Figure 5: Benchmarking GEN mechanism for $k\leq 20$ across FMs and Ts. Best NORM combination for each row used implicitly. AGG is fixed to max. CD diagram proves that k-means weight generation is significantly better than all other methods.
Initially, we limit the number of generated proxies ($k\leq 20$). Results in fig. 5 show how k-means, using as many proxies as possible (in this case, $20$) outperforms by $4\%$ on average accuracy compared to all the other GEN methods. The CD diagram illustrates its statistical significance in ranking. Furthermore, while k-medoids with $20$ proxies is computationally expensive, it is statistically on par with mean, and covariance maximization, furthest-point sampling and random selection show even weaker performances. We find similar results for $k\leq 5$, where k-means outperforms the other methods as well (see fig. 13).
Figure 6: Benchmarking different GEN methods with resnet18 on CIFAR-10. All combinations employ L2 for all NORM.
As the number of proxies ($k$) increases, k-means continues to be the best GEN method. An example for resnet18 and CIFAR-10 can be found in fig. 6. All methods converge towards the point of imprinting (saving) all data ($k=\text{all}$), even surpassing it in the case of k-means. Due to its superior performance, we mainly focus on k-means in the remainder of the analysis.
Normalization (NORM).
To investigate the role of normalization, we compare all the different NORM methods, focusing on k-means as GEN with $k=1,5,20$. For $k=1$ and varying NORMpost (while taking best values for NORMpre and NORMinf implicitly), fig. 7 shows that for weight normalization, L2 is by far the best choice. quantile and none normalization perform significantly worse.
NORMpost Avg. acc. %
L2 87.76
quantile 80.71
none 84.03
[…]
Figure 7: Benchmarking NORMpost mechanism across FMs and Ts. The best NORMpre and NORMinf combinations for each row are used implicitly. GEN is fixed to mean (that is, $k=1$) and AGG is fixed to max. The CD diagram shows the statistical significance of L2 as weight normalization NORMpost.
NORMpre NORMinf Avg. acc.
none L2 87.76
none none 88.36
L2 L2 87.72
Table 1: Benchmarking NORMpre and NORMinf mechanisms across FMs and Ts. NORMpost is fixed to L2, GEN to mean, and AGG to max. No statistically significant differences were found.
Keeping L2 for NORMpost fixed, we find no statistical differences between the different combinations of NORMpre and NORMinf. Its performances can be found in table 1. For larger values of $k$, the differences among NORMpost become even more pronounced, but for NORMpre and NORMinf, it stays statistically indifferent for L2 weight normalization (see tables 2 and 14 for all combinations at once with $k=1$, and tables 3 and 15 for $k=20$).
Henceforth, we limit all the succeeding experiments to Qi’s [4] normalization, that is, using L2 for all NORM. We choose this combination of normalizations to specifically capture cosine similarity in max aggregation.
Aggregation (AGG).
In addition to max, we study the effect of m-nn as an aggregation method. Recall that max is a special case of m-nn when $m=1$ (as NORMpost is set to L2). We investigate different values for $m\in\{1,3,5,20,50\}$.
AGG Avg. acc. %
5-nn 93.86
3-nn 93.67
20-nn 93.73
1-nn 92.97
max 92.97
50-nn 93.15
[…]
Figure 8: Benchmarking AGG mechanism across FMs and Ts. GEN is fixed to all ($k=\text{all}$), that is, imprinting (saving) all data to weights. L2 normalization is used for all NORM. The CD diagram shows statistical significance of 3-nn, 5-nn, and 20-nn over max aggregation.
When all data is imprinted, fig. 8 shows that using m-nn aggregation for $m\in\{3,5,20\}$ is slightly better than max.
AGG Avg. acc. %
1-nn 91.47
max 91.47
3-nn 90.97
5-nn 90.54
20-nn 87.57
[…]
Figure 9: Benchmarking AGG mechanism across FMs and Ts. GEN is fixed to k-means with $k=20$. L2 normalization is used for all NORM. The CD diagram shows that max is the best-performing aggregation method.
With $k=20$ and k-means as GEN, max ($=$1-nn) aggregation becomes the top performing combination (see fig. 9). Furthermore, the reduction of proxies (from all $\approx 6000$ per class to $k=20$) leads to a minor decrease in accuracy of around $2\%$.
6.2 Few-Shot Scenario
We analyze the n-shot scenario with our method, that is, k-means as GEN, L2 normalizations, and max as AGG. Furthermore, we focus on the large tasks T, that contain ten classes at once. In this scenario, due to sampling only a few examples ($n$), we average over five different seeds.
Figure 10: k-means with different values for $k$ in n-shot scenarios. $95\%$ confidence intervals are shown in shaded colors. Other variables are fixed to our previously described best method of using L2 normalizations and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples.
From the results shown in fig. 10, we find that as the number of samples $n$ increases, k-means starts outperforming mean imprinting. The usage of a higher number of proxies $k$ results in even greater performance. This shift occurs at roughly 50 samples per class for MNIST and FashionMNIST, while for CIFAR-10, $k>1$ becomes prominently better at around 200 samples per class (see fig. 16 for a focus on $10\leq n\leq 200$).
6.3 Neural Collapse and Number of Proxies
Figure 11 depicts the neural collapse measurement $\mathcal{NC}_{1}$ (see eq. 1) for the three tasks containing all of MNIST, FashionMNIST, CIFAR-10, as well as the 100 ImageNet tasks with remapped labels as explained in section 5. We can see that ImageNet has a close-to-zero $\mathcal{NC}_{1}$ score, which increases linearly when adding more classes to each label (i.e., increasing multi-modality). As for other datasets, CIFAR-10 is generally more collapsed according to its low value of $\mathcal{NC}_{1}$. We hypothesize that this is due to the similarity of its categories to those appearing in ImageNet. Apart from that, the $\mathcal{NC}_{1}$ for Transformer-based architectures is much lower and therefore they are more collapsed compared to the CNN-inspired FMs. Architectural differences are further explained in section A.2.
Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$. The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted in dots. When compared to the $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures. A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
For the same data, fig. 12 depicts accuracy over a varying number of proxies $k$ inferred from k-means. A prominent peak at $k=d$ can be inferred for every FM, and reflects that $d$ class proxies lead to the best result for $d$-modal class distributions. The fact that CIFAR-10 has the lowest $\mathcal{NC}_{1}$ (see fig. 11) is reflected by flat curves over $k$. This confirms that the $\mathcal{NC}_{1}$ score is a significant indicator of multi-modality. Namely, a higher $\mathcal{NC}_{1}$ score indicates the benefits of using a higher number of proxies.
Furthermore, increasing $k$ for the ImageNet sets has a much larger effect on the CNN-based models. We argue that this is because of their higher values of $\mathcal{NC}_{1}$ and investigate this more deeply in section A.2.
Figure 12: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN. $95\%$ confidence intervals are shown in shaded colors. L2 for all NORM and max as aggregation are used. Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines. In all four plots, peaks in accuracy at $k=d$ can be inferred. This confirms the connection between the effect of using multiple proxies and the collapse of the data.
7 Discussion
We present a new framework to analyze the three main components relevant to weight imprinting, namely, weight generation, normalization, and aggregation. Within this framework, state-of-the-art imprinting strategies become special cases. This allows for a comprehensive analysis of different approaches through systematic experiments and leads us to generalize to a new, best-performing imprinting strategy. That is, using k-means weight generation with L2 normalizations and max aggregation outperforms all previously studied methods (see fig. 4).
k-means generates better weights than mean.
In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small. Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed in [4], highlighting its advantage in few-shot scenarios.
L2 weight normalization is essential for strong performance.
The max aggregation directly scales with the magnitude of the weights. Normalization (NORMpost) ensures that all class weights contribute equally to the output. Nearest neighbor (l-nn) aggregation is not as affected by the lack of normalization, since it uses Euclidean distance. Although still part of common procedure, normalizations for embeddings (NORMpre and NORMinf) appear to have minimal impact on performance.
With max aggregation, there is no need to store all data.
While nearest neighbor (m-nn) aggregation (AGG) performs well when all data is saved (e.g., when there are no storage constraints), max aggregation with limited number of representative proxies (e.g., $k$-means) is an efficient alternative without a substantial loss in performance.
Neural collapse proves the efficacy of imprinting.
During training, the last-layer weights of a FM collapse to their respective class means. This proves the success of mean imprinting on known classes. New, out-of-distribution data, however, often shows less collapse, making it beneficial to imprint more than one proxy.
Limitations.
Since our experiments are limited to a small selection of foundation models and tasks, running additional experiments could strengthen the statistical significance. While imprinting alone provides an efficient solution to transfer learning, we do not investigate the benefit of combining it with optimization methods like gradient-based learning when more samples become available. This combination could use imprinting as initialization or apply metric learning to improve imprinting capabilities.
Future Work.
The usage of both weight and activation sparsity as in [42] could change the within- and between-class variability in favor of using a higher number of proxies. Synaptic intelligence approaches like the weight saturation presented in this work are paths of further study. We use the penultimate layer embeddings for generating the classifier weights. An interesting area of study could be extracting embeddings from previous layers of the FM for this purpose. Lately, the study in [43] showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable. Apart from that, it would be interesting to imprint the weights of other layers as well (see, for example, [10]).
8 Conclusion
We investigated imprinting as an efficient method for transfer learning based on foundation models. Within our new framework, we found a new imprinting strategy that outperforms all previously studied ones. The phenomenon of neural collapse provides theoretical proof for its success.
Acknowledgments and Disclosure of Funding
Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12, as well as the European Union under the grant project 101079894 (COMFORT - Improving Urologic Cancer Care with Artificial Intelligence Solutions). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA). Neither the European Union nor the granting authority can be held responsible for them. Furthermore, we would like to thank Viet Anh Khoa Tran for initial discussions about the neural collapse phenomenon.
Author Contributions
JW contributed to the development of the framework, conducting experiments and evaluated the findings. GA was responsible for investigating NC measures and overall contribution to the project. MK contributed to extending the framework and handling data preparation. AF provided critical feedback on the presentation of the results and contributed to refining the manuscript. AL, ER, and FG provided supervision, contributed to the overall concepts presented, and to refining the manuscript.
Appendix A Appendix
A.1 Additional Tables and Critical Difference Diagrams
We provide additional tables and critical difference (CD) diagrams that are referenced and put into context in the main paper.
GEN $k$ Avg. acc. %
k-means 5 89.57
mean 1 87.76
k-medoids 5 85.59
k-cov-max 5 82.83
k-random 5 76.41
k-fps 2 63.90
[…]
Figure 13: Benchmarking GEN mechanism for $k\leq 5$ across FMs and Ts. Best NORM combination for each row is used implicitly. AGG is fixed to max. CD diagram depicts statistical significance of k-means as GEN. See fig. 5 for $k\leq 20$.
NORMinf NORMpre NORMpost Avg. acc. %
L2 none L2 87.76
none none L2 87.76
L2 L2 L2 87.72
L2 L2 quantile 80.71
none none quantile 80.72
L2 none quantile 80.72
L2 L2 none 84.03
L2 none none 73.75
none none none 73.75
Table 2: Benchmarking NORM across FMs and Ts shows crucial effect of L2 normalization. GEN is fixed to mean and AGG to max. See fig. 14 for average ranks.
Figure 14: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 2 for further details and average accuracies.
NORMinf NORMpre NORMpost Avg. acc. %
L2 L2 L2 91.47
L2 none L2 91.45
none none L2 91.45
L2 L2 quantile 90.91
L2 L2 none 89.97
L2 none quantile 79.84
none none quantile 79.84
L2 none none 79.24
none none none 79.24
Table 3: Benchmarking NORM across FMs and Ts. GEN is fixed to k-means with $k=20$ and AGG to max. See fig. 15 for average ranks.
Figure 15: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 3 for further details and average accuracies.
Figure 16: k-means with different values for $k$ in n-shot scenarios with focus on $10\leq n\leq 200$. $95\%$ confidence intervals shown in shaded colors. Other variables are fixed to our previously described best method of using L2 for NORM and max as AGG. Note that only data for the meaningful case of $k\leq n$ is shown. It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples. See fig. 10 for more values of $n$.
A.2 Differences between Foundation Models
While an in-depth comparison of foundation models is beyond the scope of this paper, we believe it is important to highlight key observations. In particular, fig. 11 shows significantly lower $\mathcal{NC}_{1}$ scores for vit_b_16 and swin_b on their pre-training ImageNet data compared to the resnet models. We hypothesize that this difference is primarily due to model size and training regimes. The Transformer-based architectures (vit_b_16 and swin_b) have a considerably higher parameter count ($\approx 87$M) than the resnet models (11.7M and 25.6M, respectively). Additionally, vit_b_16 and swin_b were trained for three times as many epochs (300 vs. 90) while using a substantially lower learning rate (0.003 and 0.01 vs. 0.1). Notably, the embedding dimensions of these models are comparable, meaning that the observed differences in $\mathcal{NC}_{1}$ scores cannot be attributed to differences in representation dimensionality. Instead, we argue that the combination of larger model size, extended training duration, and lower learning rates likely contributes to greater overfitting, leading to more pronounced collapse.
Figure 17: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN. $95\%$ confidence intervals are shown in shaded colors. We set NORMpost and NORMinf to L2, and NORMpre to none. Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines. Besides the prominent peaks in accuracy at $k=d$ (as already observed in fig. 12), a consistent dip between $k=1$ and $k=d$ appears in Transformer-based models.
Figure 17, similar to fig. 12, illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs). The key difference in this figure is the use of none for NORMpre instead of L2. This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig. 12, where L2 is used as NORMpre, and does not occur at all in the resnet models. We hypothesize that this difference arises from the distinct embedding distributions of CNN and Transformer architectures (see, e.g., [14, Figure S2]).

### v2

@@
 We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization.
-We determine those proxies through clustering and propose a novel variant of imprinting that outperforms previous work.
+We determine proxies through clustering and propose a novel variant of imprinting that outperforms previous work.
 We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time.
 Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes.
+Finally, we publicly release our code at https://github.com/DATEXIS/multi-imprinting/.
 1 Introduction
@@
 We present a unifying framework that enables a systematic comparison of existing imprinting techniques.
-More precisely, we can generalize prior work by decomposing imprinting into three principal steps (see fig.
+More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig.
 1).
-During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (multiple proxies).
+During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (proxies).
 Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors.
@@
 The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments.
-We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization.
+We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization, outperforming previously studied methods, as depicted in fig.
+2.
 Neural Collapse.
@@
 The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$.
-The final output of the test data in T is computed by an aggregation (AGG) mechanism.
+The final output for the test data in T is computed by an aggregation (AGG) mechanism.
 Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively.
 During inference, embeddings are normalized according to NORMinf (not shown here).
+Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
+Qi et al. [4] L2 mean L2 L2 max 87.72
+Hosoda et al. [14] none mean quantile none max 80.71
+Janson et al. [15] none mean none none 1-nn 87.64
+Ours L2 k-means L2 L2 max 91.48
+[…]
+[table]A table beside a figure
+Figure 2: Previously studied imprinting strategies are special cases within our framework.
+The framework enables the creation of a novel configuration ("Ours") that outperforms previous work across FMs and Ts by a large margin with statistical significance.
+Here, $k=20$ is chosen.
 2 Related Work
@@
 Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification.
-We focus on variability collapse ($\mathcal{NC}_{1}$), using the equation 1 from [13]:
+We focus on variability collapse ($\mathcal{NC}_{1}$) as in [13]:
 $\mathcal{NC}_{1}=\tfrac{1}{c}\mathrm{trace}(\Sigma_{W}{\Sigma^{+}_{B}}),$ (1)
@@
 We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture.
-In neural collapse investigations (section 6.3), we also use resnet50 [34] and swin_b [36].
-All of them are pre-trained on ImageNet-1K (ILSVRC 2012) [37].
+In neural collapse investigations (section 6.3), we also work with resnet50 [34] and swin_b [36].
+All four models are pre-trained on ImageNet-1K (ILSVRC 2012) [37].
 To generate the embeddings, we use PyTorch’s torchvision models.
 Tasks.
-To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10  [^2] https://www.cs.toronto.edu/~kriz/cifar.html, each containing 10 classes.
+To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10 [40], each containing 10 classes.
 We mainly focus on the three T containing all ten classes.
@@
 $\{0,2,4,6,8\}$.
-This random selection of $3\cdot 4=12$ tasks adds variation to our evaluation.
+This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
 In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet).
@@
 See fig.
-3 for a simplified example.
+4 for a simplified illustration.
 We do this random sampling 10 times which results in $10\cdot 10=100$ tasks.
+Furthermore, to ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
 Figure 4: Combining multiple classes into one to create tasks with multi-modal class distributions. Example: "$d$ in 1", $d=2,3$.
 Scale.
-In total, we run approximately 350 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds.
+In total, we run approximately 150 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds.
 This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
@@
 We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [41].
-[^3] Code used to generate these diagrams can be found here [41].
+[^2] The code used to generate these diagrams is inspired by [42].
 In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy.
@@
 6.1 Best Imprinting Strategy
-Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
-Qi et al. [4] L2 mean L2 L2 max 87.72
-Hosoda et al. [14] none mean quantile none max 82.15
-Janson et al. [22] none mean none none 1-nn 87.65
-Ours L2 k-means L2 L2 max 91.47
-[…]
-[table]A table beside a figure
-Figure 4: Previously studied imprinting strategies are special cases within our framework.
-The framework enables the creation of a novel configuration ("Ours") that outperforms previous work across FMs and Ts by a large margin with statistical significance.
-Here, $k=20$ is chosen.
 We provide a comparison between memory-constrained methods used for imprinting on foundation models in fig.
-4, namely, [4, 14, 22], as well as a novel configuration that results from our framework which we call “Ours”.
+2, namely, [4, 14, 15], as well as a novel configuration that results from our framework which we call “Ours”.
 We investigate the impact of using m-nn aggregation on all data afterward.
@@
 GEN $k$ Avg. acc. %
-k-means 20 91.95
+k-means 20 91.44
 k-medoids 20 87.31
@@
 k-cov-max 20 84.74
-k-random 20 82.79
-k-fps 2 63.90
+k-random 20 82.86
+k-fps 20 66.23
 […]
@@
 CD diagram proves that k-means weight generation is significantly better than all other methods.
+Figure 6: Benchmarking different GEN methods with resnet18 on CIFAR-10. All combinations employ L2 for all NORM.
 Initially, we limit the number of generated proxies ($k\leq 20$).
@@
 13).
-Figure 6: Benchmarking different GEN methods with resnet18 on CIFAR-10. All combinations employ L2 for all NORM.
 As the number of proxies ($k$) increases, k-means continues to be the best GEN method.
@@
 L2 87.76
-quantile 80.71
-none 84.03
+quantile 80.70
+none 84.04
 […]
@@
 none L2 87.76
-none none 88.36
+none none 87.76
 L2 L2 87.72
@@
 5-nn 93.86
-3-nn 93.67
+3-nn 93.68
 20-nn 93.73
@@
 AGG Avg. acc. %
-1-nn 91.47
-max 91.47
-3-nn 90.97
-5-nn 90.54
-20-nn 87.57
+1-nn 91.48
+max 91.48
+3-nn 91.01
+5-nn 90.61
+20-nn 87.64
 […]
@@
 Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$.
-The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted in dots.
-When compared to the $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures.
+The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted as dots.
+When compared to $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures.
 A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
@@
 That is, using k-means weight generation with L2 normalizations and max aggregation outperforms all previously studied methods (see fig.
-4).
+2).
 k-means generates better weights than mean.
@@
 GEN $k$ Avg. acc. %
-k-means 5 89.57
+k-means 5 89.61
 mean 1 87.76
 k-medoids 5 85.59
-k-cov-max 5 82.83
-k-random 5 76.41
-k-fps 2 63.90
+k-cov-max 5 82.82
+k-random 5 76.40
+k-fps 5 63.74
 […]
@@
 L2 L2 L2 87.72
-L2 L2 quantile 80.71
-none none quantile 80.72
-L2 none quantile 80.72
-L2 L2 none 84.03
+L2 L2 quantile 80.70
+L2 none quantile 80.71
+none none quantile 80.71
+L2 L2 none 84.04
 L2 none none 73.75
@@
 NORMinf NORMpre NORMpost Avg. acc. %
-L2 L2 L2 91.47
-L2 none L2 91.45
-none none L2 91.45
-L2 L2 quantile 90.91
-L2 L2 none 89.97
-L2 none quantile 79.84
-none none quantile 79.84
-L2 none none 79.24
-none none none 79.24
+L2 none L2 91.44
+none none L2 91.44
+L2 L2 L2 91.48
+L2 L2 quantile 90.86
+L2 L2 none 89.99
+L2 none quantile 80.36
+none none quantile 80.36
+L2 none none 79.22
+none none none 79.22
 Table 3: Benchmarking NORM across FMs and Ts. GEN is fixed to k-means with $k=20$ and AGG to max. See fig. 15 for average ranks.

### v3

@@
 Abstract
-The capacity of a foundation model allows for adaptation to new downstream tasks.
-Weight imprinting is a universal and efficient method to fulfill this purpose.
-It has been reinvented several times, but it has not been systematically studied.
-In this paper, we propose a framework for imprinting, identifying three main components: generation, normalization, and aggregation.
-This allows us to conduct an in-depth analysis of imprinting and a comparison of the existing work.
-We reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization.
-We determine proxies through clustering and propose a novel variant of imprinting that outperforms previous work.
-We motivate this by the neural collapse phenomenon – an important connection that we can draw for the first time.
-Our results show an increase of up to 4% in challenging scenarios with complex data distributions for new classes.
-Finally, we publicly release our code at https://github.com/DATEXIS/multi-imprinting/.
+The capacity of foundation models allows for their application to new, unseen tasks.
+The adaptation to such tasks is called transfer learning.
+An efficient transfer learning method that circumvents parameter optimization is imprinting.
+The conceptual differences between studies on imprinting form the basis of our systematic investigation.
+In this work, we propose the general IMPRINT framework, identifying three main components: generation, normalization, and aggregation.
+Through the lens of this framework, we conduct an in-depth analysis and a comparison of the existing methods.
+Our findings reveal the benefits of representing novel data with multiple proxies in the generation step and show the importance of proper normalization.
+Beyond an extensive analytical grounding, our framework enables us to propose a novel variant of imprinting which outperforms previous work on transfer learning tasks by 4%.
+This variant determines proxies through clustering motivated by the neural collapse phenomenon – a connection that we draw for the first time.
+We publicly release our code at https://github.com/DATEXIS/IMPRINT.
 1 Introduction
 In machine learning applications, training models from scratch is often not viable due to limitations in data and compute.
-A popular solution is to apply transfer learning [1, 2] based on foundation models (FMs) [3] that are pre-trained on a large amount of data.
-A common approach in practice to adapt an FM to a novel task is to freeze its parameters and replace the output layer with a new head, e.g., for classification.
+A popular solution is to apply transfer learning (transferlearning3; transferlearning2) based on foundation models (FMs) (foundationmodels) that are pre-trained on a large amount of data.
+To tune a FM to a novel task, e.g., for classification, a common procedure is to freeze the model parameters and replace the output layer with a new head.
+A particularly simple method for implementing such a new head was proposed by qi-imprinting and coined imprinting.
 Imprinting.
-Qi et al.
-[4] propose a simple solution for few-shot classification, called imprinting.
-Namely, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean.
-These class means are representatives of the classes, which we generally call proxies.
-This results in an efficient method without the need for gradient-based optimization.
-A plethora of studies have emerged surveying this technique by adding complexity and adaptability [5, 6, 7, 8, 9, 10, 11].
-Despite many adaptations, imprinting lacks a systematic comparison that unifies them.
-Understanding its variations could unlock greater efficiency and performance across many fields, making the method even more versatile and impactful.
+In the original work by qi-imprinting, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean.
+These class means are representatives of the classes, which we denote as proxies.
+In general, we refer to imprinting as efficient learning methods without the need for cross-class statistics or gradient-based optimization.
+A plethora of studies have emerged surveying this technique by adding complexity and adaptability (fewshot-metagenweights; masked-imprinting; hypersphere-imprinting; imprint-multilabel; imprint-objdetec; multiple-projection-head-imprinting; yan2023few).
+Despite its many adaptations, imprinting lacks a systematic comparison that unifies them.
+Understanding its variations could unlock greater efficiency across many fields, making the method even more versatile and impactful.
+Applications.
+In particular, while in some practical applications, accuracy is prioritized over computational efficiency, the latter does become a critical requirement in scenarios where computational resources are severely constrained, e.g., in the chemical and polymer processing industries.
+Here, battery-powered edge compute is essential and frequent retraining or large-scale optimization is infeasible.
+Imprinting has proven particularly effective in these contexts of edge-embedded devices (hypersphere-imprinting).
+For instance, imprintingrobot implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting).
+Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (coralai), that allows users to add new classes from a few examples without recompiling the model.
+Very recently, (belal2025fsid) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks.
+Apart from that, ncmbaseline extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
+Figure 1: Overview of our IMPRINT framework.
+The foundation model FM is frozen and shows neural collapse.
+The weight generator (GEN) uses training data from a novel task T to consecutively generate one or more weight vectors (proxies) per class $1,\dotsc,C$ in T.
+In inference, the final output for the test data in T is computed by an aggregation (AGG) mechanism.
+Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively.
+During inference, embeddings are normalized according to NORMinf.
 Framework.
-We present a unifying framework that enables a systematic comparison of existing imprinting techniques.
-More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig.
-1).
-During generation (GEN) of weights, the network selects representative data samples and generates one or more weight vectors per class (proxies).
-Normalization (NORM) is crucial, as the network needs to balance its generated weight vectors.
+We present IMPRINT, a framework that enables a comprehensive analysis of existing imprinting techniques.
+More precisely, we generalize prior work by decomposing imprinting into three principal steps (see fig.˜1).
+During generation (GEN) of weights, the method selects representative data samples and generates one or more weight vectors (proxies) per class.
+Normalization (NORM) is crucial, as the generated weights need to be balanced.
 Aggregation (AGG) entails the computation of the final output, e.g., a class label.
-The efficiency of imprinting enables us to perform a comprehensive analysis with a large number of experiments.
-We present a new, best-performing imprinting strategy using multi-modal weight imprinting in combination with the correct way of normalization, outperforming previously studied methods, as depicted in fig.
-2.
+The computational efficiency of imprinting allows us to perform a large number of experiments.
+Through IMPRINT, we are able to propose a novel, best-performing imprinting strategy using multi-proxy weight imprinting in combination with $L^{2}$ normalization, outperforming previously studied methods, as depicted in fig.˜2.
 Neural Collapse.
-We investigate a recently discovered phenomenon called neural collapse [12], which provides a compelling explanation for why imprinting works.
-According to this phenomenon, when neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means [12, 13].
-Our investigation proves that a measurement of neural collapse provides insights about imprinting.
+When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (papyan2020prevalence; zhu2021geometric).
+We investigate this phenomenon as a potential explanation for when and why imprinting works.
+Our analysis proves that there exists a relationship between a measure of neural collapse and the success of imprinting.
+Since quantification of this phenomenon is possible in a post-hoc fashion over the FM features, we believe that these insights will contribute to further development of imprinting methods, as well as, training regimes of FMs that are more suitable for transfer learning via imprinting.
 Contributions.
 In summary, our main results and contributions are:
-  • We deconstruct weight imprinting into a framework composed of generation, normalization, and aggregation, and discuss variations for each of them, identifying prior work as special cases (section 3).
-To the best of our knowledge, we are the first to conduct a comprehensive analysis of imprinting to this scale (section 5).
-  • We present a new imprinting method utilizing $k$-means clustering for weight generation (section 6.1) and show its benefits in certain few-shot scenarios (section 6.2).
-  • To the best of our knowledge, we are the first to identify a connection between imprinting success and measures of neural collapse (section 6.3).
-We make the source code to reproduce our results publicly available. [^1] https://github.com/DATEXIS/multi-imprinting/
-Figure 1: Overview of our imprinting framework.
-The foundation model FM is frozen and shows neural collapse.
-The weight generator (GEN) uses training data from a novel task T to generate one or more weight vectors (proxies) per class $1,\dotsc,c$.
-The final output for the test data in T is computed by an aggregation (AGG) mechanism.
-Embeddings and generated weights are normalized according to NORMpre and NORMpost, respectively.
-During inference, embeddings are normalized according to NORMinf (not shown here).
-Paper NORMpre GEN NORMpost NORMinf AGG Avg. acc. %
-Qi et al. [4] L2 mean L2 L2 max 87.72
-Hosoda et al. [14] none mean quantile none max 80.71
-Janson et al. [15] none mean none none 1-nn 87.64
-Ours L2 k-means L2 L2 max 91.48
+  • We deconstruct weight imprinting into the IMPRINT framework composed of generation, normalization, and aggregation, and discuss variations for each of them, while identifying prior work as special cases (section˜3.1).
+To the best of our knowledge, we are the first to conduct a comprehensive analysis of imprinting to this scale (section˜4).
+  • We present a new, outperforming imprinting method utilizing $k$-means clustering for weight generation (section˜5.1) and show its benefits in certain low-data regimes (section˜5.2).
+  • As far as we are aware, we are the first to identify a relationship between the degree of neural collapse and imprinting success (section˜5.3).
 […]
-[table]A table beside a figure
-Figure 2: Previously studied imprinting strategies are special cases within our framework.
-The framework enables the creation of a novel configuration ("Ours") that outperforms previous work across FMs and Ts by a large margin with statistical significance.
-Here, $k=20$ is chosen.
+[…]
+Figure 2: Previously studied imprinting strategies are special cases within IMPRINT.
+We evaluate 12 different classification tasks Ts derived from MNIST, FashionMNIST, and CIFAR-10, each with 10 classes or subsets thereof, and 4 pre-trained models FMs (resnet18, resnet50, vit_b_16, swin_b).
+The proposed configuration (“Ours”) derived from IMPRINT outperforms previous work across FMs and Ts by a large margin with statistical significance, as confirmed by the critical difference (CD) diagram below.
+Since absolute accuracies vary substantially across tasks and models (as reflected by the large standard deviations (std)), this rank-based aggregation is used for fair comparison.
+Here, $k=20$ is used, highlighting the gain of using multiple proxies per class.
+For reference, the gray row reports an oracle method that uses cross-class feature statistics to generate weights (see section˜A.5).
+It is not an imprinting method and therefore not directly comparable to the imprinting-based approaches above.
+Nonetheless, the results indicate that our method substantially narrows the gap between single-proxy mean imprinting and this oracle baseline.
 2 Related Work
-Imprinting and Few-Shot Learning.
-Weight imprinting was introduced in [4] for the few-shot learning scenario.
-It is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples.
-Qi et al.
-[4] find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms.
-However, we are not limiting the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
-Imprinting has also been applied to object detection [5, 6], multi-label classification [8], semantic segmentation [10], and in combination with an attention mechanism to generate weights for the novel classes in a few-shot classification task [11].
-Hosoda et al.
-[14] apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights.
+Imprinting and Low-Data Regimes.
+Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by qi-imprinting for the few-shot learning scenario.
+The authors find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms.
+In contract, in our work we do not limit the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
+Imprinting has been applied to object detection (imprint-objdetec; yan2023few), multi-label classification (imprint-multilabel), semantic segmentation (masked-imprinting), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (fewshot-metagenweights).
+itsdone apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights.
 We consider this as one normalization scheme in our framework.
-Zhang et al.
-[16] apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available.
+covidimprint apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available.
 They speculate whether normalization is a constraint in their imprinting model.
-Before the era of deep learning, Mensink et al.
-[17] analyze the transferability of hand-crafted image features.
-They use a "nearest class multiple centroids" (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm.
+Before the era of deep learning, first-nearestclassmean analyze the transferability of hand-crafted image features.
+They use a “nearest class multiple centroids” (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm.
 In combination with metric learning, they compare favorably against the $m$-nearest neighbor algorithm.
@@
 Transfer Learning.
-Our work is related to the use of embedding vectors extracted with pre-trained models, which is one of the most straightforward transfer learning techniques, since the seminal works in computer vision [18] and natural language processing [19].
-Kornblith et al.
-[20] showed that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks.
-In addition, Huh et al.
-[21] provided insights into the required quality of pre-training data.
+Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (donahue_decaf_2014) and natural language processing (devlin_bert_2019).
+kornblith_better_2019 show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks.
+In addition, huh_what_2016 provide insights into the required quality of pre-training data.
 Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
 Continual Learning (CL).
-Class means have also been used as proxies in CL.
 Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL.
-Rebuffi et al.
-[22] dynamically select a subset of examples for each class and update internal representations via gradient descent.
+icarl dynamically select a subset of examples for each class and update internal representations via gradient descent.
 They use a nearest mean classifier (NMC) with respect to the saved examples.
-Janson et al.
-[15] use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings.
+ncmbaseline use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings.
 However, they do not investigate the effect of normalization and using multiple proxies.
-Findings of [23] show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored.
+Findings of imprintallknn show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored.
 In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
 Neural Collapse (NC).
-The phenomenon of NC was identified by [12] and refers to the convergence of the last-layer weight vectors to class means.
-It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) [13, 24, 25], but complete neural collapse is practically unrealistic [26].
-In transfer learning, Galanti et al.
-[27] show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios.
+The phenomenon of NC was identified by papyan2020prevalence and refers to the convergence of the last-layer weight vectors to class means.
+It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (zhu2021geometric; neuralcollapse2; neuralcollapse3).
+Nevertheless, complete neural collapse is practically unrealistic (DBLP:conf/icml/TirerHN23).
+In transfer learning, DBLP:conf/iclr/Galanti0H22 show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios.
 In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
-3 Imprinting Framework
-In order to find out how to best set the classifier weights of a foundational model in downstream tasks T, we create a framework (see fig.
-1) that encompasses many different combinations, all of which work without gradient-based training.
-Thereby, we can unify all the existing imprinting strategies described in section 2.
-We analyze multi-class classification scenarios in that we do not separate into base and new classes, but focus on all classes in a novel T at the same time.
-To investigate the effect of the number of samples given, we look at n-shot ($n\in\mathbb{N}$) scenarios.
-For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the regime of few-shot learning.
+3 Methods
+3.1 IMPRINT
+In order to find out how to best and efficiently set the classifier weights of a foundational model FM in new, previously unseen tasks T, we create the IMPRINT framework (see fig.˜1) that generalizes previous work, specifically all the methods that work without access to cross-class statistics and gradient-based training.
+Thereby, we unify all the existing imprinting strategies described in section˜2.
 Overview.
 We analyze the effect of weight generation (GEN), normalizations ($\texttt{NORM}=\{\texttt{NORM}\textsubscript{{pre}},\texttt{NORM}\textsubscript{{post}},\texttt{NORM}\textsubscript{{inf}}\}$), and aggregation (AGG).
-The framework depicted in fig.
-1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted.
+The IMPRINT framework depicted in fig.˜1 consists of three main building blocks: a foundation model FM, a weight generator GEN, and extendable classifier weights that are imprinted.
 The FM remains frozen throughout the experiments.
@@
 Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre.
-The generated weight vectors per class are called proxies, prototypes, or representatives [28, 29, 30, 31].
+The generated weight vectors per class are referred to as proxies, prototypes, or representatives (metric-proxies; prototypical; convolutional-prototypes; semanticdriftCL).
 These proxies are normalized according to NORMpost.
-As in [4], we do not use bias values.
+As in the work by qi-imprinting, we do not use bias values.
 To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
-Special Imprinting Cases.
-Previously proposed imprinting methods can be defined as a special case of our framework.
-Figure 2 in section 6.1 provides an overview of existing imprinting strategies listed in literature for foundational models and benchmarks them with the new best-performing one we find through our framework.
-In total, we inspect all possible combinations (including variations in models, tasks, and seeds).
+Generalization of Previous Methods.
+Previously proposed imprinting methods can be defined as special cases of our framework.
+We design IMPRINT such that every method can be defined by a single combination of GEN, NORM, and AGG, of which we inspect all possible combinations.
 Weight Generation (GEN).
 The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights.
-In contrast to [4] which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in [17].
-This enables non-linear classification.
-We denote the number of proxies as $k$, ranging between $1$ and the number of samples.
-We investigate the following operations conducted per class to generate its proxies:
+In contrast to qi-imprinting which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in first-nearestclassmean and enable non-linear classification.
+We denote the number of proxies per class as $k$, ranging between $1$ and the number of samples, and investigate the following operations conducted per class to generate those:
   • all: All embeddings (denoted as $k=\text{all}$).
@@
   • mean: The mean of all embeddings.
-  • k-means: $k$-means cluster centers. $k=1$ is the same as mean.
-  • k-medoids: $k$-medoids cluster centers.
-  • k-cov-max (covariance-maximization): Top $k$ embeddings by covariance.
-  • k-fps (farthest-point sampling): Iteratively selecting $k$ embeddings, such that it maximizes the distance from already selected ones (starting with random sample).
+  • k-means: $k$-means cluster centers using KMeans from sklearn (scikit-learn). $k=1$ is the same as mean.
+  • k-medoids: $k$-medoids cluster centers using KMedoids from sklearn.
+  • k-cov-max (covariance-maximization): Top $k$ embeddings by covariance (in code: proxies = embeddings[torch.argsort(torch.sum(torch.cov(embeddings), dim=0), descending=True)[:k]]).
+  • k-fps (farthest-point sampling): Iteratively selecting $k$ embeddings, such that it maximizes the distance from already selected ones (starting with a random sample).
 We choose this diverse list of methods to cover a wide range of approaches, ranging from heuristics (e.g., k-fps) to more complex algorithms (e.g., k-means).
-Note that only mean and k-means generate proxies that are not included in the given samples.
+Note that only mean and k-means generate proxies beyond existing samples by producing synthetic cluster centers in embedding space.
+In comparison, k-medoids must choose actual samples as cluster centers (analogous to a median).
+None of these methods use cross-class statistics or gradient-based optimization.
+We also note an analogy to associative memory models, interpreting imprinting as a memory update process following a covariance rule, as detailed in section˜A.3.
 Normalization (NORM).
@@
 L2 normalization can be applied to embeddings before GEN via NORMpre, to the generated weights via NORMpost, and to embeddings in inference via NORMinf.
-In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean $\|\cdot\|_{2}$ length.
-quantile normalization [32, 33] can only be applied to generated weights.
+In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean length $\|\cdot\|_{2}$.
+quantile normalization (quantnorm1; quantnorm2) can only be applied to generated weights.
 This non-linear operation distributes weights equally.
-Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class, and the reference distribution changes accordingly.
+Recall that if more than one class is contained in T ($c>1$), GEN is performed for each class consecutively, and the reference distribution changes accordingly.
 In particular, for the first class there is no reference distribution to map to.
-This is different from [14], where new weights are matched to the distribution of the original classifier weights of the FM.
+This is different from itsdone, where new weights are matched to the distribution of the original classifier weights of the FM.
 Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
 Aggregation (AGG).
-There are various ways to use the generated weights per class during inference, especially when $k>1$.
-We focus on two different modes, max and m-nn.
+There are various ways to use the generated weights (proxies) per class during inference, especially when $k>1$.
+We focus on two different modes, max and $m$-nearest neighbor (m-nn).
 The former, max, computes the inner product of the input embedding and the imprinted weights and outputs the class label with the maximum activation.
 The latter, m-nn, uses the class weights as keys and the embeddings as values, and chooses the final winning output class via the $m$-nearest neighbor algorithm.
-The m-nn voting is weighted by the inverse of the distances to their nearest neighbor, turning it into weighted majority voting.
-Note that max is the same as 1-nn in the case of L2 for NORMpost, since for any fixed embedding vector $v$ and variable proxy $w$, the argmin of $\|v-w\|^{2}=\|v\|^{2}-2\langle v,w\rangle+\|w\|^{2}$, calculated by 1-nn, is the same as the argmax of the inner product $\langle v,w\rangle$ calculated in max.
-4 Measurement of Neural Collapse
-Neural collapse (NC) [12] refers to the phenomenon that occurs on the last-layer classifier weights of neural networks in the terminal phase of training (TPT).
-When the network is trained well beyond zero training error, the learned embeddings of each class, assuming balanced classes, collapse to their class means.
+The m-nn voting is weighted by the inverse of the Euclidean distances to their nearest neighbor, turning it into weighted majority voting.
+We also experimented with alternative distance functions (Cosine, Manhattan) and uniform weighting, but found differences well within statistical noise.
+Using the Mahalanobis distance proved computationally prohibitive.
+More elaborate voting mechanisms, such as learned top-$k$ attention or entropy-based filtering, are conceivable, but extend well beyond the minimalist imprinting paradigm and therefore left for future work.
+Note that max is the same as 1-nn in the case of L2 for NORMpost, since for any fixed embedding vector $\boldsymbol{\mathbf{v}}$ and variable proxy $\boldsymbol{\mathbf{w}}$, the argmin of $\|\boldsymbol{\mathbf{v}}-\boldsymbol{\mathbf{w}}\|^{2}=\|\boldsymbol{\mathbf{v}}\|^{2}-2\langle\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{w}}\rangle+\|\boldsymbol{\mathbf{w}}\|^{2}$, calculated by 1-nn, is the same as the argmax of the inner product $\langle\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{w}}\rangle$ calculated in max.
+3.2 Quantifying Neural Collapse
+Neural collapse (NC) (papyan2020prevalence) characterizes the state of the features produced by a classification neural network after training to near zero training loss.
+Namely, the learned embeddings of each class converge, i.e., collapse, to their class means.
 These globally centered class means and classifier weights form a simplex equiangular tight frame (ETF) – a collection of equal length and maximally equiangular vectors, that maximize the between-class variability.
 This results in an optimal linearly separable state for classification.
-In fig.
-3 (left), we illustrate the collapse of a FM on its pre-training data.
+In fig.˜3 (left), we illustrate the collapse of a FM on its pre-training data.
 The newly arrived data T from a different dataset is distributed more unevenly across the embedding space (right).
-Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification.
-We focus on variability collapse ($\mathcal{NC}_{1}$) as in [13]:
-$\mathcal{NC}_{1}=\tfrac{1}{c}\mathrm{trace}(\Sigma_{W}{\Sigma^{+}_{B}}),$ (1)
-where $\Sigma_{W},\Sigma_{B}\in\mathbb{R}^{l\times l}$ are within- and between-class covariance matrices, respectively, $l$ is the dimension of the embedding vector, $c$ is the number of classes, and + symbolizes the pseudo-inverse.
-Based on the equation, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse.
-In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig.
-11).
-Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $c$.
-According to NC, imprinting the mean, as originally done in [4], is best when $\mathcal{NC}_{1}$ is small.
-We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class.
-We investigate this in section 6.3.
 Figure 3: Left: The embeddings of the pre-training data, after being used to train the foundation model FM, show neural collapse, as each class ($o_{1},\dotsc,o_{4}$) is evenly separated in space and accumulates around their respective class means.
 Right: For a novel task with classes $c_{1},c_{2}$ (pink and brown) scatter around the collapsed pre-trained classes (gray).
+Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification.
+We focus on variability collapse ($\mathcal{NC}_{1}$) as in zhu2021geometric.
+Given a foundation model FM and a finite dataset containing $C$ classes with (for simplicity) $N$ samples per class, we use its $l$-dimensional embeddings $\{\boldsymbol{\mathbf{h}}_{c,i}\}_{1\leq c\leq C,\;1\leq i\leq N}$ given by FM to define the global feature mean $\boldsymbol{\mathbf{h}}_{G}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, class means $\overline{\boldsymbol{\mathbf{h}}}_{c}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, the within-class covariance matrix $\mathbf{\Sigma}_{W}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)^{\top}$, and the between-class covariance matrix $\mathbf{\Sigma}_{B}\coloneqq\frac{1}{C}\sum_{c=1}^{C}\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)^{\top}$ to finally compute
+[…]
+where + symbolizes the pseudo-inverse.
+Based on eq.˜1, an $\mathcal{NC}_{1}$ score closer to zero signifies a higher collapse.
+In contrast, an increase in multi-modality of data leads to a higher $\mathcal{NC}_{1}$ score (as analyzed in fig.˜11(a)).
+Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $C$.
+According to NC, imprinting the mean, as originally done by qi-imprinting, is best when $\mathcal{NC}_{1}$ is small.
+We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class.
+We investigate this in section˜5.3.
+3.3 Significance Testing with Critical Difference Diagrams
+Our experiments compare large numbers of imprinting configurations across tasks with substantially different achievable accuracy levels.
+To obtain scale-independent statistical comparisons, we base our analysis on ranking (dis-)agreements rather than raw accuracy.
+Non-parametric rank-based tests are well suited for this setting.
+Following cddiag, we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric.
+Empirical evidence in cddiag further suggests that these tests have good power in classifier comparison scenarios.
+Concretely, for each evaluation instance (i.e., foundation model FM and task T) we compute the classification accuracy of all configurations and assign ranks (rank 1 = highest accuracy).
+This yields a matrix of ranks with one row per configuration and one column per evaluation instance.
+On this matrix we first apply the Friedman test to assess whether there are any overall differences between configurations.
+If the Friedman test rejects the global null hypothesis at significance level $\alpha=0.05$, we proceed with a post-hoc, pairwise analysis.
+Specifically, for every pair of configurations we run a two-sided Wilcoxon signed-rank test across evaluation instances.
+The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (holm1979simple).
+Critical difference (CD) diagrams summarize both the average performance and the significance structure.
+The horizontal axis shows the average rank of each configuration across all evaluation instances, with lower ranks indicating better performance.
+Two configurations are connected by a thick horizontal line if their Wilcoxon–Holm comparison does not show a significant difference.
 4 Experimental Setup
-Models.
-We use resnet18 [34] and vit_b_16 [35] as FMs, one CNN-based and one Transformer-based architecture.
-In neural collapse investigations (section 6.3), we also work with resnet50 [34] and swin_b [36].
-All four models are pre-trained on ImageNet-1K (ILSVRC 2012) [37].
+Foundation Models FM.
+We use resnet18, resnet50 (resnet), vit_b_16 (vit), and swin_b (swin) as FMs, two CNN-based and two Transformer-based architectures.
+All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (imagenet).
 To generate the embeddings, we use PyTorch’s torchvision models.
-Tasks.
-To find out the best imprinting strategy within our framework, we focus on tasks T created from the datasets MNIST [38], FashionMNIST [39], and CIFAR-10 [40], each containing 10 classes.
+Tasks T.
+We analyze multi-class classification scenarios without separating base and new classes, instead focusing on all classes within a novel T at the same time.
+To investigate the effect of the number of samples given, we look at n-shot ($n\in\mathbb{N}$) scenarios.
+For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the low-data regime.
+To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (mnist), FashionMNIST (fashionmnist), and CIFAR-10 (cifar10), each containing 10 classes.
 We mainly focus on the three T containing all ten classes.
@@
 This random selection of $3\cdot 4=12$ tasks adds variation to our evaluations.
-In the investigations of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet).
+Neural Collapse.
+In the analysis of neural collapse (NC), we also look at the FMs’ pre-training data (ImageNet).
 As its test set is not available, we use its validation set in $\mathcal{NC}_{1}$ computations.
-For ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis.
-These tasks are called "$d$ in $1$", $d=1,\dotsc,10$, each containing $10$ different labels.
+Furthermore, for ImageNet, we relabel data by combining multiple classes into one label to simulate multi-modal class distributions for an in-depth NC analysis.
+These tasks are called $d$ in 11, $d=1,\dotsc,10$, each containing $10$ different labels.
 More precisely, we take 100 random classes from ImageNet and sequentially map the first $d$ to label $1$, the second $d$ to label $2$, etc., until we reach 10 distinct labels.
-See fig.
-4 for a simplified illustration.
-We do this random sampling 10 times which results in $10\cdot 10=100$ tasks.
-Furthermore, to ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
-Figure 4: Combining multiple classes into one to create tasks with multi-modal class distributions. Example: "$d$ in 1", $d=2,3$.
+See fig.˜4 for a simplified illustration.
+Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (mnistm), SVHN (svhn), and USPS (usps).
+The resulting dataset, which we refer to as CombiDigits, exhibits reduced collapse due to the greater distributional diversity within each class.
+To ensure scale invariance in covariance-based NC measurements, all embeddings are $L^{2}$-normalized before computing $\mathcal{NC}_{1}$.
 Scale.
-In total, we run approximately 150 000 experiments, varying across the imprinting components, foundation models, tasks, and seeds.
-This is feasible with minimal effort as imprinting is a highly efficient method that operates without relying on gradient descent or other non-linear optimization techniques.
+In total, we ran approximately 500 000 experiments, varying imprinting components, foundation models, tasks, and seeds.
+This is feasible with minimal effort as imprinting is a highly efficient method: core steps such as weight generation (GEN), normalization (NORM), and aggregation (AGG) are linear in dataset size or number of proxies and parallelize efficiently within each step.
+Details on experimental infrastructure and parallelization are provided in section˜A.6.
+Figure 4: Combining multiple classes into one to create tasks with multi-modal class distributions. Simplified example for $d$ in 1”, $d=1,2,3$, with only six (instead of 100) original classes.
 Evaluation.
-Throughout our experiments, the median accuracy on the test set for three different seeds is reported, if not otherwise specified.
-In sections 6.1 and 6.2, we investigate the imprinting performance by varying the FM (2) and T (12).
-We then sort the combinations by their final accuracy.
-There are $2\cdot 12=24$ potentially different ranks for each of the combinations.
-We show the average rank, average accuracy, and statistical significance in ranking (dis-)agreements through critical difference (CD) diagrams as presented in [41].
-[^2] The code used to generate these diagrams is inspired by [42].
-In the CD diagrams, a thick horizontal line indicates a group of combinations that are not significantly different from each other in terms of accuracy.
-We consider differences significant if $p<0.05$.
-In experiments with neural collapse (section 6.3), we investigate four FMs on 100 ImageNet tasks and the three tasks containing all of MNIST, FashionMNIST, and CIFAR-10, respectively.
+Unless stated otherwise, we report the median test accuracy over three different seeds.
+In sections˜5.1 and 5.2, we evaluate the imprinting performance by varying the FM (4) and T (12) and report average accuracy and standard deviation (std) across these.
+Due to the heterogeneity of models and tasks, large std are expected.
+Therefore, overall method comparisons are based on ranks rather than absolute accuracies.
+Methods are ranked by their median accuracy, yielding $4\cdot 12=48$ potentially different ranks.
+We report the average rank and assess statistical significance of ranking (dis-)agreements using critical difference (CD) diagrams as explained in section˜3.3.
+The code used to generate these diagrams is inspired by cddiag-code.
+In experiments with neural collapse (section˜5.3), we investigate the four FMs on 100 random ImageNet tasks with remapped labels (as explained above) and the four tasks containing all of MNIST, FashionMNIST, CIFAR-10, and CombiDigits, respectively.
 5 Results
-Our main experimental insights are:
-  1. 1.
-    Our imprinting framework generalizes previous methods, and we find a new superior imprinting strategy (section 6.1).
-  2. 2.
-    We show that our strategy is beneficial in few-shot scenarios with as little as 50 samples per class (section 6.2).
-  3. 3.
-    We identify a correlation between imprinting success utilizing multiple proxies and measures of neural collapse (section 6.3).
+Our main experimental insights are: 1.
+Our IMPRINT framework generalizes previous methods, and we find a new superior imprinting strategy (section˜5.1).
+2.
+We show that our novel strategy is even beneficial in low-data regimes with as little as 50 samples per class (section˜5.2).
+3.
+We identify a correlation between imprinting success utilizing multiple proxies and a measure of neural collapse (section˜5.3).
 5.1 Best Imprinting Strategy
-We provide a comparison between memory-constrained methods used for imprinting on foundation models in fig.
-2, namely, [4, 14, 15], as well as a novel configuration that results from our framework which we call “Ours”.
-We investigate the impact of using m-nn aggregation on all data afterward.
-We focus on $k=20$ and find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all other approaches by a margin of $4\%$ with statistical significance, as can be seen in the CD diagram.
-Below, we analyze each of the components of the framework separately.
-Weight Generation (GEN).
-To analyze the impact of GEN, we first focus on the max aggregation.
-For the weight generation analysis, we do not fix NORM, but simply show the run with the best NORM combination, if not otherwise specified.
-The m-nn aggregation and different values for NORM are analyzed later in this section.
-GEN $k$ Avg. acc. %
-k-means 20 91.44
-k-medoids 20 87.31
-mean 1 87.76
-k-cov-max 20 84.74
-k-random 20 82.86
-k-fps 20 66.23
+We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by qi-imprinting; itsdone; ncmbaseline, as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT.
+Focusing on $k=20$, we find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all previously studied approaches by a margin of $4\%$ on average with statistical significance.
+For reference, we additionally report an oracle baseline that uses cross-class feature statistics to generate its weights, representing an upper bound that is not constrained by imprinting.
+The results indicate that our new imprinting method significantly narrows the gap between single-proxy mean imprinting and this oracle baseline.
+The impact of using the unconstrained m-nn aggregation on all data is investigated at the end of this section. Next, we analyze each of the components of IMPRINT separately.
 […]
+[table]A table beside a figure GEN $k$ Avg.
+acc.
+% $\pm$ std k-means 20 91.04 $\pm 6.26$ k-medoids 20 87.01 $\pm 8.23$ mean 1 86.84 $\pm 7.80$ k-cov-max 20 83.98 $\pm 9.56$ k-random 20 82.14 $\pm 10.15$ k-fps 20 65.56 $\pm 12.32$ Refer to caption
 Figure 5: Benchmarking GEN mechanism for $k\leq 20$ across FMs and Ts.
@@
 CD diagram proves that k-means weight generation is significantly better than all other methods.
-Figure 6: Benchmarking different GEN methods with resnet18 on CIFAR-10. All combinations employ L2 for all NORM.
-Initially, we limit the number of generated proxies ($k\leq 20$).
-Results in fig.
-5 show how k-means, using as many proxies as possible (in this case, $20$) outperforms by $4\%$ on average accuracy compared to all the other GEN methods.
+Weight Generation (GEN).
+To assess the impact of GEN, we first focus on the max aggregation and do not fix NORM, but simply show the run with the best NORM combination, if not otherwise specified.
+The m-nn aggregation and different values for NORM are analyzed later in this section.
+Initially, we limit the number of generated proxies to $k\leq 20$.
+Results in fig.˜5 show how k-means, using as many proxies as possible (in this case, $20$) outperforms by $4\%$ on average accuracy compared to all the other GEN methods.
 The CD diagram illustrates its statistical significance in ranking.
-Furthermore, while k-medoids with $20$ proxies is computationally expensive, it is statistically on par with mean, and covariance maximization, furthest-point sampling and random selection show even weaker performances.
-We find similar results for $k\leq 5$, where k-means outperforms the other methods as well (see fig.
-13).
-As the number of proxies ($k$) increases, k-means continues to be the best GEN method.
-An example for resnet18 and CIFAR-10 can be found in fig.
-6.
+Furthermore, while k-medoids with $20$ proxies (which necessarily have to be among the given samples, see section˜3.1) is computationally expensive, it is on par with mean, and covariance maximization, furthest-point sampling and random selection show even weaker performances.
+We find similar results for $k\leq 5$, where k-means outperforms the other methods as well (see fig.˜A.1).
+Figure 6: Benchmarking different GEN methods with resnet18 on all of CIFAR-10 shows superiority of k-means proxies. All combinations employ L2 for all NORM and max as AGG.
+As the number of proxies $k$ increases, k-means continues to be the best GEN method.
+An example for resnet18 and CIFAR-10 can be found in fig.˜6.
 All methods converge towards the point of imprinting (saving) all data ($k=\text{all}$), even surpassing it in the case of k-means.
 Due to its superior performance, we mainly focus on k-means in the remainder of the analysis.
+Learned Weights.
+Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by harun2025good, which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here.
+We extend this method by combining it with k-means, forming k-least-squares, and observe in section˜A.5 that using multiple proxies per class instead of a single weight vector can still improve performance, particularly on datasets with high $\mathcal{NC}_{1}$.
 Normalization (NORM).
-To investigate the role of normalization, we compare all the different NORM methods, focusing on k-means as GEN with $k=1,5,20$.
-For $k=1$ and varying NORMpost (while taking best values for NORMpre and NORMinf implicitly), fig.
-7 shows that for weight normalization, L2 is by far the best choice.
-quantile and none normalization perform significantly worse.
-NORMpost Avg. acc. %
-L2 87.76
-quantile 80.70
-none 84.04
+We compare all the different NORM methods, focusing on k-means as GEN.
+For $k=1$ and varying NORMpost (while taking best values for NORMpre and NORMinf implicitly), fig.˜7 shows that L2 is the best choice for weight normalization.
+quantile and none normalization both perform worse.
 […]
+[table]A table beside a figure NORMpost Avg. acc. % $\pm$ std L2 86.84 $\pm 7.80$ quantile 82.90 $\pm 12.87$ none 83.26 $\pm 9.18$ Refer to caption
 Figure 7: Benchmarking NORMpost mechanism across FMs and Ts.
@@
 GEN is fixed to mean (that is, $k=1$) and AGG is fixed to max.
-The CD diagram shows the statistical significance of L2 as weight normalization NORMpost.
-NORMpre NORMinf Avg. acc.
-none L2 87.76
-none none 87.76
-L2 L2 87.72
+The CD diagram shows the statistical significance of L2 as the best weight normalization NORMpost.
 Table 1: Benchmarking NORMpre and NORMinf mechanisms across FMs and Ts. NORMpost is fixed to L2, GEN to mean, and AGG to max. No statistically significant differences were found.
-Keeping L2 for NORMpost fixed, we find no statistical differences between the different combinations of NORMpre and NORMinf.
-Its performances can be found in table 1.
-For larger values of $k$, the differences among NORMpost become even more pronounced, but for NORMpre and NORMinf, it stays statistically indifferent for L2 weight normalization (see tables 2 and 14 for all combinations at once with $k=1$, and tables 3 and 15 for $k=20$).
-Henceforth, we limit all the succeeding experiments to Qi’s [4] normalization, that is, using L2 for all NORM.
-We choose this combination of normalizations to specifically capture cosine similarity in max aggregation.
-Aggregation (AGG).
-In addition to max, we study the effect of m-nn as an aggregation method.
-Recall that max is a special case of m-nn when $m=1$ (as NORMpost is set to L2).
-We investigate different values for $m\in\{1,3,5,20,50\}$.
-AGG Avg. acc. %
-5-nn 93.86
-3-nn 93.68
-20-nn 93.73
-1-nn 92.97
-max 92.97
-50-nn 93.15
+NORMpre NORMinf Avg. acc. % $\pm$ std
+none L2 86.84 $\pm 7.80$
+none none 86.84 $\pm 7.80$
+L2 L2 86.79 $\pm 7.83$
+Keeping L2 for NORMpost fixed, we find no statistical differences between the different combinations of NORMpre and NORMinf (see table˜1).
+For larger values of $k$, the differences among NORMpost become even more pronounced.
+However, the performance of NORMpre and NORMinf remains statistically indifferent for L2 weight normalization (see fig.˜A.2 for all combinations at once with $k=1$, and fig.˜A.3 for $k=20$).
+We restrict all the subsequent experiments to using L2 across all NORM following qi-imprinting.
+This combination of normalizations is chosen to specifically model cosine similarity within max aggregation.
 […]
+[table]A table beside a figure AGG Avg.
+acc.
+% $\pm$ std 5-$n$ 93.74 $\pm 4.97$ 3-$n$ 93.50 $\pm 5.22$ 20-$n$ 93.56 $\pm 4.93$ 1-$n$ 92.81 $\pm 5.69$ max 92.81 $\pm 5.69$ 50-$n$ 92.91 $\pm 5.25$ Refer to caption
 Figure 8: Benchmarking AGG mechanism across FMs and Ts.
@@
 L2 normalization is used for all NORM.
-The CD diagram shows statistical significance of 3-nn, 5-nn, and 20-nn over max aggregation.
-When all data is imprinted, fig. 8 shows that using m-nn aggregation for $m\in\{3,5,20\}$ is slightly better than max.
-AGG Avg. acc. %
-1-nn 91.48
-max 91.48
-3-nn 91.01
-5-nn 90.61
-20-nn 87.64
+The CD diagram shows statistical significance of 3-$n$, 5-$n$, and 20-$n$ over max aggregation.
 […]
+[table]A table beside a figure AGG Avg. acc. % $\pm$ std 1-$n$ 91.06 $\pm 6.21$ max 91.06 $\pm 6.21$ 3-$n$ 90.59 $\pm 6.16$ 5-$n$ 90.12 $\pm 6.21$ 20-$n$ 87.05 $\pm 7.46$ Refer to caption
 Figure 9: Benchmarking AGG mechanism across FMs and Ts.
@@
 The CD diagram shows that max is the best-performing aggregation method.
-With $k=20$ and k-means as GEN, max ($=$1-nn) aggregation becomes the top performing combination (see fig.
-9).
-Furthermore, the reduction of proxies (from all $\approx 6000$ per class to $k=20$) leads to a minor decrease in accuracy of around $2\%$.
-6.2 Few-Shot Scenario
-We analyze the n-shot scenario with our method, that is, k-means as GEN, L2 normalizations, and max as AGG.
-Furthermore, we focus on the large tasks T, that contain ten classes at once.
-In this scenario, due to sampling only a few examples ($n$), we average over five different seeds.
+Aggregation (AGG).
+In addition to max, we study the effect of using $m$-nearest neighbor (m-nn) as an aggregation method.
+Recall that max is a special case of m-nn when $m=1$ (as NORMpost is set to L2).
+We investigate different values for $m\in\{1,3,5,20,50\}$.
+When all data is imprinted, fig.˜8 shows that using m-nn aggregation for $m\in\{3,5,20\}$ is slightly better than max.
+With $k=20$ and k-means as GEN, max (==1-$n$) aggregation becomes the top performing combination (see fig.˜9).
+Furthermore, the reduction of proxies (from all ($\approx 6000$ per class) to $k=20$) leads to a decrease in accuracy of less than $3\%$.
 Figure 10: k-means with different values for $k$ in n-shot scenarios.
-$95\%$ confidence intervals are shown in shaded colors.
+Shaded areas indicate $95\%$ confidence intervals (CIs) over five different seeds.
 Other variables are fixed to our previously described best method of using L2 normalizations and max as AGG.
 Note that only data for the meaningful case of $k\leq n$ is shown.
-It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples.
-From the results shown in fig.
-10, we find that as the number of samples $n$ increases, k-means starts outperforming mean imprinting.
-The usage of a higher number of proxies $k$ results in even greater performance.
-This shift occurs at roughly 50 samples per class for MNIST and FashionMNIST, while for CIFAR-10, $k>1$ becomes prominently better at around 200 samples per class (see fig.
-16 for a focus on $10\leq n\leq 400$).
+For MNIST and FashionMNIST, mean ceases to be the best strategy already at only roughly 50 samples.
+5.2 Low-Data Regime
+We analyze the n-shot scenario with our method (k-means as GEN, L2 as NORM, and max as AGG).
+Furthermore, we focus on the large tasks T containing all ten classes of MNIST, FashionMNIST, resp.
+CIFAR-10 at once.
+In this scenario, due to sampling only a few examples $n$, we average over five (instead of three) different seeds.
+From the results in fig.˜10, we find that as the number of samples $n$ increases, k-means starts to outperform mean imprinting.
+The use of more proxies $k$ further amplifies this performance gain.
+The shift occurs at roughly 50 samples per class for MNIST and FashionMNIST, while for CIFAR-10, $k>1$ becomes prominently better at around 200 samples per class (see fig.˜A.4 for a display of all FMs focused on $10\leq n\leq 400$).
 5.3 Neural Collapse and Number of Proxies
-Figure 11 depicts the neural collapse measurement $\mathcal{NC}_{1}$ (see eq.
-1) for the three tasks containing all of MNIST, FashionMNIST, CIFAR-10, as well as the 100 ImageNet tasks with remapped labels as explained in section 5.
-We can see that ImageNet has a close-to-zero $\mathcal{NC}_{1}$ score, which increases linearly when adding more classes to each label (i.e., increasing multi-modality).
-As for other datasets, CIFAR-10 is generally more collapsed according to its low value of $\mathcal{NC}_{1}$.
+Figure˜11(a) depicts the neural collapse measurement $\mathcal{NC}_{1}$ (see eq.˜1) for the 100 random ImageNet tasks with remapped labels as explained in section˜4, as well as the four tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp.
+CombiDigits.
+It can be inferred that ImageNet has a close-to-zero $\mathcal{NC}_{1}$ score, which increases linearly when adding more classes to each label (i.e., increasing multi-modality).
+As for the other datasets, CIFAR-10 is generally more collapsed according to its low value of $\mathcal{NC}_{1}$, which even falls below $1$ for the Transformer-based FMs.
 We hypothesize that this is due to the similarity of its categories to those appearing in ImageNet.
-Apart from that, the $\mathcal{NC}_{1}$ for Transformer-based architectures is much lower and therefore they are more collapsed compared to the CNN-inspired FMs.
-Architectural differences are further explained in section A.2.
-Figure 11: Average $\mathcal{NC}_{1}$ and $95\%$ confidence intervals of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$.
-The $\mathcal{NC}_{1}$ for the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are depicted as dots.
-When compared to $\mathcal{NC}_{1}$ on the pre-training set (ImageNet), these sets are less collapsed on the Transformer-based architectures.
-A clear linear relationship across $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
-For the same data, fig.
-12 depicts accuracy over a varying number of proxies $k$ inferred from k-means.
-A prominent peak at $k=d$ can be inferred for every FM, and reflects that $d$ class proxies lead to the best result for $d$-modal class distributions.
-The fact that CIFAR-10 has the lowest $\mathcal{NC}_{1}$ (see fig.
-11) is reflected by flat curves over $k$.
-This confirms that the $\mathcal{NC}_{1}$ score is a significant indicator of multi-modality.
-Namely, a higher $\mathcal{NC}_{1}$ score indicates the benefits of using a higher number of proxies.
-Furthermore, increasing $k$ for the ImageNet sets has a much larger effect on the CNN-based models.
-We argue that this is because of their higher values of $\mathcal{NC}_{1}$ and investigate this more deeply in section A.2.
-Figure 12: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN.
-$95\%$ confidence intervals are shown in shaded colors.
-L2 for all NORM and max as aggregation are used.
-Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines.
-In all four plots, peaks in accuracy at $k=d$ can be inferred.
+By design, the synthetic CombiDigits dataset, introduced in section˜4, has a very high $\mathcal{NC}_{1}$.
+Apart from that, $\mathcal{NC}_{1}$ of the ImageNet data for the Transformer-based architectures are much lower and therefore they are more collapsed compared to the CNN-based FMs.
+These architectural differences are further investigated in section˜A.4.
+(a) A clear linear relationship between the neural collapse measure $\mathcal{NC}_{1}$ and $d$ can be inferred for all FMs, i.e., increased multi-modality implies less collapse.
+The $\mathcal{NC}_{1}$ of other datasets (CombiDigits in particular) is much higher across all FMs, while only vit_b_16 and swin_b get an $\mathcal{NC}_{1}$ of less than one on CIFAR-10.
+(b) Accuracy over number of proxies $k$ used in k-means together with L2 for all NORM and max as aggregation.
+In all four subplots, peaks in accuracy at $k=d$ can be inferred.
 This confirms the connection between the effect of using multiple proxies and the collapse of the data.
-7 Discussion
-We present a new framework to analyze the three main components relevant to weight imprinting, namely, weight generation, normalization, and aggregation.
-Within this framework, state-of-the-art imprinting strategies become special cases.
-This allows for a comprehensive analysis of different approaches through systematic experiments and leads us to generalize to a new, best-performing imprinting strategy.
-That is, using k-means weight generation with L2 normalizations and max aggregation outperforms all previously studied methods (see fig.
-2).
+Figure 11: Averaged $\mathcal{NC}_{1}$ resp.
+accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$.
+Shaded areas indicate $95\%$ CIs over three different seeds.
+Values for the tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp.
+CombiDigits at once are shown in dotted styles.
+For the same data, fig.˜11(b) depicts accuracy over a varying number of proxies $k$ inferred from k-means.
+A prominent peak at $k=d$ can be inferred for every FM and reflects that $d$ class proxies lead to the best result for $d$-modal class distributions.
+Furthermore, increasing $k$ for the ImageNet sets has a much larger effect on the CNN-based FMs.
+We argue that this is because of their higher values of $\mathcal{NC}_{1}$, indicating less neural collapse.
+In section˜A.4, we analyze architectural differences and training setups to explain these variations.
+The fact that CIFAR-10 has the lowest $\mathcal{NC}_{1}$ scores (see fig.˜11(a)) is reflected by flat green curves over $k$.
+Figure˜12 renders this connection more precise: For datasets with large $\mathcal{NC}_{1}$, i.e., where intra-class variability surpasses inter-class variability, using more than a single proxy per class ($k>1$) yields a clear performance gain over mean imprinting ($k=1$).
+Figure 12: Accuracy (%) gain when using $k=5$ instead of $k=1$ proxies per class with k-means, plotted against the neural collapse measure $\mathcal{NC}_{1}$ (log-scaled axis).
+Colors indicate different FMs.
+Outlined markers correspond to non-ImageNet datasets (MNIST, FashionMNIST, CIFAR-10, and CombiDigits), while filled markers denote averaged results from our random ImageNet label remappings.
+The dashed gray line shows a least-squares fit ($R^{2}=0.36$), summarizing an approximate log-linear dependence between $\mathcal{NC}_{1}$ and the gains from using $k>1$ proxies.
+Across all points, the Spearman rank correlation is $\rho=0.82$ ($p<0.0001$), indicating a statistically significant positive association.
+Consistent with this trend, for regimes with $\mathcal{NC}_{1}>1$ single-proxy mean imprinting ($k=1$) appears to become substantially suboptimal compared to using multiple proxies ($k>1$).
+6 Conclusion
+We present a new framework, IMPRINT, to analyze three main components relevant to weight imprinting, namely, weight generation, normalization, and aggregation.
+Within IMPRINT, state-of-the-art imprinting strategies become special cases.
+This allows for a comprehensive analysis of different approaches through systematic experiments and leads us to generalize to a new, best-performing imprinting variant.
+That is, using k-means weight generation with L2 normalizations and max aggregation, which outperforms all previously studied methods (see fig.˜2).
 k-means generates better weights than mean.
 In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small.
-Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed in [4], highlighting its advantage in few-shot scenarios.
+Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by qi-imprinting, highlighting its advantage in low-data regimes.
 L2 weight normalization is essential for strong performance.
 The max aggregation directly scales with the magnitude of the weights.
-Normalization (NORMpost) ensures that all class weights contribute equally to the output.
+Normalization (NORMpost) ensures that all proxies have the same magnitude, preventing differences in vector norms from disproportionately affecting classifier predictions.
 Nearest neighbor (l-nn) aggregation is not as affected by the lack of normalization, since it uses Euclidean distance.
@@
 With max aggregation, there is no need to store all data.
-While nearest neighbor (m-nn) aggregation (AGG) performs well when all data is saved (e.g., when there are no storage constraints), max aggregation with limited number of representative proxies (e.g., $k$-means) is an efficient alternative without a substantial loss in performance.
+While nearest neighbor (m-nn) aggregation (AGG) performs well when all data is saved (e.g., when there are no storage constraints), max aggregation with limited number of representative proxies (e.g., through $k$-means) is an efficient alternative without a substantial loss in performance.
 Neural collapse proves the efficacy of imprinting.
-During training, the last-layer weights of a FM collapse to their respective class means.
+During training, the last-layer weights of a FM tend to collapse to their respective class means.
 This proves the success of mean imprinting on known classes.
 New, out-of-distribution data, however, often shows less collapse, making it beneficial to imprint more than one proxy.
+Our results show that the advantage of using multiple proxies is strongly coupled to the degree of neural collapse: the accuracy gain from using $k>1$ proxies increases approximately log-linearly with $\mathcal{NC}_{1}$.
+In particular, multi-proxy imprinting yields substantial and predictable improvements on datasets with $\mathcal{NC}_{1}$ exceeding 1.
+While from a practical perspective, choosing the exact number of proxies $k$ based on pure greedy search with validation data (or as part of any AutoML pipeline) is still a valid option, our analysis provides insights into the underlying mechanism, turning $\mathcal{NC}_{1}$ into an indicator for when multi-proxy imprinting is highly benefitial.
 Limitations.
-Since our experiments are limited to a small selection of foundation models and tasks, running additional experiments could strengthen the statistical significance.
-While imprinting alone provides an efficient solution to transfer learning, we do not investigate the benefit of combining it with optimization methods like gradient-based learning when more samples become available.
-This combination could use imprinting as initialization or apply metric learning to improve imprinting capabilities.
+Our experiments are limited to foundation models for image classification and do not cover other data modalities such as audio, text, or sensor inputs.
+While IMPRINT is agnostic to modality, empirical validation of our results outside of vision is needed.
+Although imprinting alone provides an efficient solution to transfer learning, when compared to purely gradient-based learning, a gap still remains.
+We do not investigate the benefit of combining it with optimization methods such as gradient-based learning, and the choice of $k$ still requires heuristic or empirical selection rather than direct prediction.
 Future Work.
-The usage of both weight and activation sparsity as in [43] could change the within- and between-class variability in favor of using a higher number of proxies.
-Synaptic intelligence approaches like the weight saturation presented in this work are paths of further study.
-We use the penultimate layer embeddings for generating the classifier weights.
-An interesting area of study could be extracting embeddings from previous layers of the FM for this purpose.
-Lately, the study in [44] showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable.
-Apart from that, it would be interesting to imprint the weights of other layers as well (see, for example, [10]).
-8 Conclusion
-We investigated imprinting as an efficient method for transfer learning based on foundation models.
-Within our new framework, we found a new imprinting strategy that outperforms all previously studied ones.
-The phenomenon of neural collapse provides theoretical proof for its success.
-Acknowledgments and Disclosure of Funding
-Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Project-ID 528483508 - FIP 12, as well as the European Union under the grant project 101079894 (COMFORT - Improving Urologic Cancer Care with Artificial Intelligence Solutions).
+The usage of both weight and activation sparsity as shen2023reducing could change the within- and between-class variability in favor of using a higher number of proxies.
+Besides only using the penultimate layer embeddings for generating the classifier weights, an interesting area of study could be extracting embeddings from previous layers of the FM for this purpose.
+Recently, the study by marczak2025revisiting showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable.
+Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (masked-imprinting).
+Acknowledgments
+Our work is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – FIP-12 – Project-ID 528483508, as well as the European Union under the grant project 101079894 (COMFORT - Improving Urologic Cancer Care with Artificial Intelligence Solutions).
 Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or European Health and Digital Executive Agency (HADEA).
 Neither the European Union nor the granting authority can be held responsible for them.
-Furthermore, we would like to thank Viet Anh Khoa Tran for initial discussions about the neural collapse phenomenon.
+We thank Viet Anh Khoa Tran for initial discussions about the neural collapse phenomenon.
+We further thank the reviewers and the Action Editor for their detailed and constructive feedback, which substantially improved this work.
 Author Contributions
@@
 Appendix A Appendix
-A.1 Additional Tables and Critical Difference Diagrams
-We provide additional tables and critical difference (CD) diagrams that are referenced and put into context in the main paper.
-GEN $k$ Avg. acc. %
-k-means 5 89.61
-mean 1 87.76
-k-medoids 5 85.59
-k-cov-max 5 82.82
-k-random 5 76.40
-k-fps 5 63.74
+A.1 Additional Results
+We provide additional tables and critical difference (CD) diagrams that expand on the results of section˜5.
 […]
-Figure 13: Benchmarking GEN mechanism for $k\leq 5$ across FMs and Ts.
+[table]A table beside a figure GEN $k$ Avg.
+acc.
+% $\pm$ std k-means 5 89.15 $\pm 6.96$ mean 1 86.84 $\pm 7.80$ k-medoids 5 84.87 $\pm 8.76$ k-cov-max 5 82.11 $\pm 10.39$ k-random 5 75.93 $\pm 11.92$ k-fps 5 63.64 $\pm 12.29$ Refer to caption
+Figure A.1: Benchmarking GEN mechanism for $k\leq 5$ across FMs and Ts.
 Best NORM combination for each row is used implicitly.
@@
 CD diagram depicts statistical significance of k-means as GEN.
-See fig.
-5 for $k\leq 20$.
-NORMinf NORMpre NORMpost Avg. acc. %
-L2 none L2 87.76
-none none L2 87.76
-L2 L2 L2 87.72
-L2 L2 quantile 80.70
-L2 none quantile 80.71
-none none quantile 80.71
-L2 L2 none 84.04
-L2 none none 73.75
-none none none 73.75
-Table 2: Benchmarking NORM across FMs and Ts shows crucial effect of L2 normalization. GEN is fixed to mean and AGG to max. See fig. 14 for average ranks.
-Figure 14: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 2 for further details and average accuracies.
-NORMinf NORMpre NORMpost Avg. acc. %
-L2 none L2 91.44
-none none L2 91.44
-L2 L2 L2 91.48
-L2 L2 quantile 90.86
-L2 L2 none 89.99
-L2 none quantile 80.36
-none none quantile 80.36
-L2 none none 79.22
-none none none 79.22
-Table 3: Benchmarking NORM across FMs and Ts. GEN is fixed to k-means with $k=20$ and AGG to max. See fig. 15 for average ranks.
-Figure 15: CD diagram depicting statistical significance of L2 for NORMpost. Combinations are listed as "NORMinf & NORMpre & NORMpost". See table 3 for further details and average accuracies.
-Figure 16: k-means with different values for $k$ in n-shot scenarios with focus on $10\leq n\leq 400$.
-$95\%$ confidence intervals shown in shaded colors.
+See fig.˜5 for $k\leq 20$.
+[…]
+[table]A table beside a figure NORMinf NORMpre NORMpost Avg.
+acc.
+% $\pm$ std L2 none L2 86.84 $\pm 7.80$ none none L2 86.84 $\pm 7.80$ L2 L2 L2 86.79 $\pm 7.83$ L2 none quantile 82.90 $\pm 12.87$ none none quantile 82.90 $\pm 12.87$ L2 L2 quantile 82.83 $\pm 12.86$ L2 L2 none 83.26 $\pm 9.18$ L2 none none 67.66 $\pm 22.12$ none none none 67.66 $\pm 22.12$ Refer to caption
+Figure A.2: Benchmarking NORM across FMs and Ts shows crucial effect of L2 normalization.
+GEN is fixed to mean and AGG to max.
+CD diagram depicting statistical significance of L2 for NORMpost.
+Combinations are listed as NORMinf & NORMpre & NORMpost.
+[…]
+[table]A table beside a figure NORMinf NORMpre NORMpost Avg.
+acc.
+% $\pm$ std L2 none L2 91.04 $\pm 6.26$ none none L2 91.04 $\pm 6.26$ L2 L2 L2 91.06 $\pm 6.21$ L2 L2 quantile 90.51 $\pm 6.40$ L2 L2 none 89.55 $\pm 6.70$ L2 none quantile 79.21 $\pm 15.20$ none none quantile 79.21 $\pm 15.20$ L2 none none 73.53 $\pm 21.05$ none none none 73.53 $\pm 21.05$ Refer to caption
+Figure A.3: Benchmarking NORM across FMs and Ts.
+GEN is fixed to k-means with $k=20$ and AGG to max.
+CD diagram depicting statistical significance of L2 for NORMpost.
+Combinations are listed as NORMinf & NORMpre & NORMpost.”
+Figure A.4: k-means with different values for $k$ in n-shot scenarios with focus on $10\leq n\leq 400$.
+Shaded areas indicate $95\%$ confidence intervals (CIs) over five different seeds.
 Other variables are fixed to our previously described best method of using L2 for NORM and max as AGG.
 Note that only data for the meaningful case of $k\leq n$ is shown.
-It can be inferred that for MNIST and FashionMNIST, mean is not the best strategy anymore starting at roughly 50 samples.
-See fig.
-10 for more values of $n$.
-A.2 Differences between Foundation Models
-While an in-depth comparison of foundation models is beyond the scope of this paper, we believe it is important to highlight key observations.
-In particular, fig.
-11 shows significantly lower $\mathcal{NC}_{1}$ scores for vit_b_16 and swin_b on their pre-training ImageNet data compared to the resnet models.
+For MNIST and FashionMNIST, mean ceases to be the best strategy already at only roughly 50 samples.
+See fig.˜10 for more values of $n$.
+A.2 Datasets
+We briefly describe the datasets used in our experiments.
+ImageNet (imagenet).
+We use the ILSVRC 2012 version (commonly called ImageNet-1K) containing 1 000 classes and 1.2M training images.
+Since the test set is not publicly available, we use the validation set (50 000 images) as a stand-in.
+For neural collapse investigations, we construct tasks by randomly grouping $d$ classes into one label, producing “$d$ in 1” tasks as explained in section˜4.
+MNIST (mnist).
+A benchmark dataset of handwritten digits (0–9), consisting of 60 000 training and 10 000 test grayscale images of size $28\times 28$.
+FashionMNIST (fashionmnist).
+A drop-in replacement for MNIST with the same format and number of samples, containing grayscale images of fashion items across 10 classes.
+CIFAR-10 (cifar10).
+A dataset of $32\times 32$ RGB images covering 10 object classes, with 50 000 training and 10 000 test samples.
+MNIST-M (mnistm), SVHN (svhn), USPS (usps).
+Digit classification datasets each containing digits 0-9 with domain-specific visual characteristics.
+MNIST-M applies color and texture transformations to MNIST digits, yielding 60 000 training and 10 000 RGB images of size $28\times 28$.
+SVHN consists of digit crops from house numbers in Google Street View, totaling 73 257 training and 26 032 test images of $32\times 32$ in RGB.
+USPS contains scanned and normalized handwritten digits in $16\times 16$ grayscale, split into 7 219 training and 2 007 test images.
+CombiDigits (Ours).
+A synthetic dataset constructed by merging all classes from MNIST, MNIST-M, SVHN, and USPS.
+Each class label corresponds to a digit (0–9) and fig.˜A.5 shows example images of class “6”.
+The dataset includes significant visual heterogeneity across sources and thus simulates multi-modal, non-collapsed class distributions.
+Figure A.5: Twenty sample images from class “6” of the CombiDigits dataset.
+Each row corresponds to one of the four source datasets: MNIST, MNIST-M, SVHN, and USPS (top to bottom), illustrating the intra-class heterogeneity across domains.
+A.3 Imprinting as Memory
+We revisit the classical idea of Bidirectional Associative Memories (BAMs) (kohonen; anderson72; nakano; anderson77) and the associated update rule (sejnowski; dayan) for storing key-value pairs as in
+$\mathbf{W}\leftarrow\mathbf{W}+\boldsymbol{\mathbf{v}}\boldsymbol{\mathbf{k}}^{\top},$
+where $\mathbf{W}\in\mathbb{R}^{l\times l}$ is a matrix and $\boldsymbol{\mathbf{k}},\boldsymbol{\mathbf{v}}\in\mathbb{R}^{d}$ are key and value vectors, respectively, to be stored therein.
+Throughout this work, imprinting can be interpreted as inserting such key-value associations into the classifier weights.
+More precisely, the generation (GEN) component extends the linear classification head by setting $\boldsymbol{\mathbf{v}}$ as a one-hot vector representing the class, and $\boldsymbol{\mathbf{k}}$ as the corresponding proxy previously computed.
+While we focus on this simple linear setting, imprinting is not limited to it, as discussed in section˜2.
+Notably, this form of direct memory update has seen renewed attention in modern architectures beyond standard query-key-value attention.
+In particular, the xLSTM model (xlstm) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
+A.4 Differences between Foundation Models
+While an in-depth comparison of foundation models is beyond the scope of this work, we believe it is important to highlight key observations we have made.
+In particular, fig.˜11(a) shows significantly lower $\mathcal{NC}_{1}$ scores for vit_b_16 and swin_b on their pre-training ImageNet data compared to the resnet models.
 We hypothesize that this difference is primarily due to model size and training regimes.
 The Transformer-based architectures (vit_b_16 and swin_b) have a considerably higher parameter count ($\approx 87$M) than the resnet models (11.7M and 25.6M, respectively).
-Additionally, vit_b_16 and swin_b were trained for three times as many epochs (300 vs.
+Additionally, vit_b_16 and swin_b were trained for more than three times as many epochs (300 vs.
 90) while using a substantially lower learning rate (0.003 and 0.01 vs.
@@
 Instead, we argue that the combination of larger model size, extended training duration, and lower learning rates likely contributes to greater overfitting, leading to more pronounced collapse.
-Figure 17: Averaged accuracy of ten random ImageNet label remappings ("$d$ in $1$") for every $d=1,\dotsc,10$ over number of proxies $k$ used for k-means in GEN.
-$95\%$ confidence intervals are shown in shaded colors.
-We set NORMpost and NORMinf to L2, and NORMpre to none.
-Accuracies of the tasks containing all of MNIST, FashionMNIST, and CIFAR-10 at once are shown in dotted lines.
-Besides the prominent peaks in accuracy at $k=d$ (as already observed in fig.
-12), a consistent dip between $k=1$ and $k=d$ appears in Transformer-based models.
-Figure 17, similar to fig.
-12, illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs).
+As discussed in section˜5.3, this enables the Transformer-based FMs to handle the ImageNet tasks with remapped labels more effectively and to achieve significantly better performance on the similarly distributed CIFAR-10 dataset.
+(a) k-means used as GEN.
+NORMpost and NORMinf are set to L2, and NORMpre to none.
+Besides the prominent peaks in accuracy at $k=d$ (as already observed in fig.˜11(b)), a consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, which was not to be seen with NORMpre set to L2 as well.
+In section˜A.4 we hypothesize that this is due to the distinct embeddings distributions of CNN- and Transformer-based architectures.
+(b) k-least-squares (see section˜A.5) used as GEN and all NORM are set to none.
+Here, the CNN-based FMs exhibit a clear accuracy peak at $k=d$, with a consistent drop for intermediate values $1<k<d$.
+For the Transformer-based FMs, however, accuracy is typically highest at $k=1$ and only improves slightly around $k=d$ but does not show a pronounced peak.
+For all FMs, performance declines with $k>d$, indicating that excessive proxy splitting harms generalization.
+This degradation does not occur in k-means GEN, which remains robust even with large $k$ (see figs.˜11(b) and 6(a)).
+Figure A.6: Averaged accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$ over number of proxies $k$.
+Shaded areas indicate $95\%$ CIs over three different seeds.
+Values for the tasks containing all of MNIST, FashionMNIST, CIFAR-10, resp.
+CombiDigits at once are shown in dotted lines.
+Figure˜6(a), similar to fig.˜11(b), illustrates the impact of varying the number of proxies on imprinting accuracy across different foundation models (FMs).
 The key difference in this figure is the use of none for NORMpre instead of L2.
-This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.
-12, where L2 is used as NORMpre, and does not occur at all in the resnet models.
-We hypothesize that this difference arises from the distinct embedding distributions of CNN and Transformer architectures (see, e.g., [14, Figure S2]).
+This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.˜11(b), where L2 is used as NORMpre, and does not occur at all in the resnet models.
+We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., itsdone).
+A.5 Learned Weights (with Multiple Proxies) and Comparisons
+harun2025good recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions.
+In contrast, our work centers on imprinting-based approaches, which avoid using gradient-based optimization and cross-class statistics, operating instead on a per-class basis with immediate availability of data.
+Nonetheless, we include the least-squares method as a supervised oracle baseline – explicitly not an imprinting method – fine-tuned for classification accuracy.
+We evaluate the performance of least-squares on the same tasks as defined in section˜4.
+Additionally, we introduce a multi-proxy extension, k-least-squares, which combines least squares with k-means clustering.
+As in the case of weight generation GEN in imprinting (where k-means outperforms mean), we find that using multiple proxies per class improves performance in settings with high $\mathcal{NC}_{1}$ as well, indicating that the benefits of multi-prototype representations persist even in this non-imprinting, supervised context.
+All experiments use no normalization (none as NORM), based on ablations confirming that additional normalization impairs performance.
+This matches expectations, since least squares outputs are already calibrated, and normalization distorts them.
+Least Squares Weights.
+For all data contained in task T, define $\mathbf{H}\in\mathbb{R}^{l\times N}$ as the collection of all feature vectors of the $N$ training samples in T obtained by applying a fixed FM, i.e., $\boldsymbol{\mathbf{h}}_{c,i}\in\mathbb{R}^{l}$ is the feature vector of the $i$-th sample in the $c$-th class.
+Recall from section˜3.2 that the within-class covariance matrix $\mathbf{\Sigma}_{W}$ is given as $\mathbb{E}_{c,i}[(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\mathbf{h}}_{c})(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\mathbf{h}}_{c})^{\top}]$ and additionally define the total covariance matrix $\mathbf{\Sigma}_{T}$ and the class-means matrix $\mathbf{M}$ as
+$\mathbf{M}=[\overline{\boldsymbol{\mathbf{h}}}_{1},\ldots,\overline{\boldsymbol{\mathbf{h}}}_{C}]\in\mathbb{R}^{l\times C},\qquad\mathbf{\Sigma}_{T}=\mathbb{E}_{c,i}[(\boldsymbol{\mathbf{h}}_{c,i}-\boldsymbol{\mathbf{h}}_{G})(\boldsymbol{\mathbf{h}}_{c,i}-\boldsymbol{\mathbf{h}}_{G})^{\top}]\in\mathbb{R}^{l\times l}.$
+From these feature statistics, we can obtain the least-squares weights via
+$\mathbf{W}_{LS}=\frac{1}{C}\mathbf{M}^{\top}(\mathbf{\Sigma}_{T}+\boldsymbol{\mathbf{h}}_{G}\boldsymbol{\mathbf{h}}_{G}^{\top}+\lambda\mathbf{I})^{-1}.$ (A.1)
+Here, $\mathbf{I}$ is the identity matrix and $\lambda$ is the weight decay.
+We set $\lambda$ to match the value used during the original training of each respective model.
+Table˜A.1 lists the specific $\lambda$ values used for all models considered in our experiments.
+Table A.1: Weight decay values $\lambda$ used for each model.
+Model $\lambda$
+resnet18, resnet50 (resnet) 0.0001
+vit_b_16 (vit) 0.1
+swin_b (swin) 0.05
+Equation˜A.1 reveals that least-squares shares structural similarities with mean imprinting, but applies a more sophisticated normalization scheme that includes both scaling and rotation.
+Crucially, it relies on cross-class statistics computed across the entire dataset, and thus is incompatible with imprinting scenarios, which operate by directly constructing class weights from the data of a single class without supervision or access to other classes.
+Algorithm 1 k-least-squares
+1:Class data $\{\mathbf{H}_{c}\}_{c=1}^{C}$, number of proxies $k$
+2:Weights $\{\mathbf{W}_{c}\}_{c=1}^{C}$ with shape $[k,l]$ per class
+3:if $k=1$ then
+4:  return standard least squares weights $\mathbf{W}_{LS}$ for each class (see eq.˜A.1)
+5:end if
+6:for each class $c$ do
+7:  Cluster $\mathbf{H}_{c}$ into $k$ clusters via $k$-means
+8:  Assign each cluster $j$ to proxy class $(c,j)$
+9:  Collect proxy samples $\{\mathbf{H}_{(c,j)}\}_{j=1}^{k}$
+10:end for
+11:Compute least squares weights $\boldsymbol{\mathbf{w}}_{(c,j)}$ jointly for all proxy classes $(c,j)$
+12:Assemble $\mathbf{W}_{c}=[\boldsymbol{\mathbf{w}}_{(c,1)},\ldots,\boldsymbol{\mathbf{w}}_{(c,k)}]$ for each original class $c$
+13:return $\{\mathbf{W}_{c}\}_{c=1}^{C}$
+Combining k-means and least-squares into k-least-squares.
+To allow for multiple proxies per class, we propose k-least-squares, a generalization of standard least squares (least-squares) that integrates clustering.
+Instead of computing a single weight vector per class using all class samples, we first partition each class’s feature set $\mathbf{H}_{c}$ into $k$ clusters using $k$-means.
+Each cluster is treated as a separate proxy class, effectively expanding the classification task from $C$ to $k\cdot C$ proxy classes.
+We then solve a single regularized least squares problem over all proxy classes at once, assigning a distinct target vector to each proxy.
+The resulting weights are grouped by their original class to yield $k$ weight vectors per class.
+The complete procedure is given in algorithm˜1.
+The results in fig.˜6(b) extend previous findings from mean and k-means for the CNN-based FMs, as accuracy increases with increasing proxy count $k$, peaks at $k=d$, and reveals a distinct dip between $k=1$ and $k=d$, confirming the weakness of using only one proxy per class.
+For Transformer-based FMs, by contrast, accuracy is often highest at $k=1$, improves slightly near $k=d$, but lacks a pronounced peak.
+Beyond $k>d$, accuracy generally declines for k-least-squares, indicating diminishing returns from excessive proxy splitting.
+In contrast, k-means GEN maintains stable performance even as $k$ grows, demonstrating robustness to increasing proxy counts.
+Comparison with k-means.
+We now compare the performance of k-means and k-least-squares as proxy generation methods, noting again that k-least-squares is not an imprinting scheme as it does not operate on a class-by-class basis as the imprinting methods presented in section˜3.1 do.
+Upon proper comparison of the best performing k-means and k-least-squares configurations, as depicted in fig.˜A.7, we observe that on our synthetic ImageNet tasks, k-least-squares does not consistently offer a substantial improvement.
+In fact, it is only on resnet50, in a highly multi-modal setting ($d=10$), that k-least-squares reaches better accuracy.
+For the Transformer-based models, k-means even performs better in these scenarios.
+While the Transformer-based models generally achieve superior performance overall (note the different y-axis scales for CNN- and Transformer-based models in fig.˜A.7), a consistent trend across all models is observed: with increasing multi-modality (increasing $d$), using more than one proxy, whether through k-means or k-least-squares, begins to outperform the single-proxy methods (mean and least-squares).
+Still, it is striking to see how effective single weights with least-squares can be.
+However, it is also noteworthy how easily k-means can bridge this gap, particularly considering that least-squares is not an actual imprinting scheme, while k-means is.
+Figure A.7: Averaged accuracy of ten random ImageNet label remappings ($d$ in 11) for every $d=1,\dotsc,10$ for mean and least-squares, and optimal k-means and k-least-squares ($k\in\{0,\dotsc,15\}$) weights.
+Shaded areas indicate $95\%$ CIs over three different seeds.
+Note the differing y-axis scales for CNN-based (resnet18, resnet50) and Transformer-based (vit_b_16, swin_b) models.
+The figure illustrates that with increasing multi-modality $d$, multi-proxy methods (k-means, k-least-squares) generally outperform single-proxy methods (mean, least-squares), and that the k-means imprinting scheme is competitive with least-squares approaches.
+Furthermore, we evaluate the k-least-squares weight generation method (using none for all NORM) across the 48 tasks proposed in the main part of our paper (see section˜4).
+least-squares reaches an average accuracy of 94.5494.54%, while 20-least-squares drops to 91.2091.20%.
+The observed decrease in accuracy can be explained by the dynamics shown in fig.˜6(b).
+In comparison to the numbers presented in fig.˜2, this shows that our $k$-means imprinting scheme significantly bridges the gap between single-proxy imprinting ($k=1$) and the optimal least squares weights.
+A.6 Computational Efficiency
+Clustering-based imprinting.
+Our k-means implementation in GEN uses sklearn.cluster.KMeans with its default parameters.
+The cost is $\mathcal{O}(Nklt)$ for $N$ samples, $k$ clusters, feature dimension $l$, and $t$ Lloyd iterations.
+While $k$-means is not computationally negligible, both assignment and update steps in each iteration parallelize naturally, and convergence typically requires only a few iterations.
+Similarly, the other steps in our imprinting framework (NORM, AGG) scale linearly with dataset size or proxy count and parallelize efficiently, resulting in no practical scalability bottlenecks even for larger or more complex datasets.
+Gradient-based optimization.
+The closed-form least-squares costs $\mathcal{O}(Nl^{2}+l^{3}+NlC)$ from the covariance computation and matrix inversion.
+If solved through stochastic gradient descent, the cost is $\mathcal{O}(NlCt)$ with $t$ epochs.
+Empirical runtime.
+Table˜A.2 shows that least-squares is faster than k-means in practice, likely because it processes all data in a single closed-form step rather than iterating class-by-class.
+Nevertheless, it should be clarified again that least-squares represents the analytic, non-iterative optimum derived from gradient minimization and lacks sequential class handling – precisely where imprinting excels (e.g., in continual or edge-learning scenarios).
+Table A.2: Average runtime in seconds for different GEN variants across all 48 tasks (see section˜4).
+To ensure transparency, all timing measurements were obtained on identical hardware (8 vCPUs on Intel Xeon Gold 6438Y+ nodes (2 sockets, 64 physical cores/128 threads)).
+Cores were not pinned to processes, which may introduce minor variance; however, we observed less than 10% variation across seeds.
+GEN variant Runtime (s)
+mean 0.0029
+5-means 1.2249
+20-means 1.7321
+least-squares 0.1593
+5-least-squares 1.3724
+20-least-squares 2.3717

### v4

@@
 In machine learning applications, training models from scratch is often not viable due to limitations in data and compute.
-A popular solution is to apply transfer learning (transferlearning3; transferlearning2) based on foundation models (FMs) (foundationmodels) that are pre-trained on a large amount of data.
+A popular solution is to apply transfer learning (Bengio, 2012; Yosinski et al., 2014) based on foundation models (FMs) (Bommasani et al., 2021) that are pre-trained on a large amount of data.
 To tune a FM to a novel task, e.g., for classification, a common procedure is to freeze the model parameters and replace the output layer with a new head.
-A particularly simple method for implementing such a new head was proposed by qi-imprinting and coined imprinting.
+A particularly simple method for implementing such a new head was proposed by Qi et al.
+(2018) and coined imprinting.
 Imprinting.
-In the original work by qi-imprinting, the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean.
+In the original work by Qi et al.
+(2018), the last-layer weight vector of a novel class is set to the normalized average of its scaled embedding vectors, i.e., its class mean.
 These class means are representatives of the classes, which we denote as proxies.
 In general, we refer to imprinting as efficient learning methods without the need for cross-class statistics or gradient-based optimization.
-A plethora of studies have emerged surveying this technique by adding complexity and adaptability (fewshot-metagenweights; masked-imprinting; hypersphere-imprinting; imprint-multilabel; imprint-objdetec; multiple-projection-head-imprinting; yan2023few).
+A plethora of studies have emerged surveying this technique by adding complexity and adaptability (Gidaris and Komodakis, 2018; Siam et al., 2019; Passalis et al., 2020; Khan et al., 2021; Li et al., 2021; Cristovao et al., 2022; Yan et al., 2023).
 Despite its many adaptations, imprinting lacks a systematic comparison that unifies them.
@@
 Here, battery-powered edge compute is essential and frequent retraining or large-scale optimization is infeasible.
-Imprinting has proven particularly effective in these contexts of edge-embedded devices (hypersphere-imprinting).
-For instance, imprintingrobot implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting).
-Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (coralai), that allows users to add new classes from a few examples without recompiling the model.
-Very recently, (belal2025fsid) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks.
-Apart from that, ncmbaseline extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
+Imprinting has proven particularly effective in these contexts of edge-embedded devices (Passalis et al., 2020).
+For instance, Zhu et al.
+(2022) implement a vision-based robotic force grasping with a variable-stiffness gripper that can safely handle both fragile and heavy objects by rapidly adapting to novel categories without retraining (Continual Learning (CL) setting).
+Industrial adoption also underscores this trend with Google’s Coral Edge TPU including an ImprintingEngine API (Coral, ), that allows users to add new classes from a few examples without recompiling the model.
+Very recently, (Belal et al., 2025) apply imprinting to spectrogram embeddings from IMU/EMG gait data, achieving efficient classification in low-data Human Activity Recognition (HAR) tasks.
+Apart from that, Janson et al.
+(2022) extend imprinting to CL, establishing a simple baseline with competitive performance compared to more sophisticated state-of-the-art CL algorithms.
 Figure 1: Overview of our IMPRINT framework.
@@
 Neural Collapse.
-When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (papyan2020prevalence; zhu2021geometric).
+When neural networks are trained to reach near-zero loss, their penultimate-layer embeddings collapse to the class means (Papyan et al., 2020; Zhu et al., 2021).
 We investigate this phenomenon as a potential explanation for when and why imprinting works.
@@
 Imprinting and Low-Data Regimes.
-Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by qi-imprinting for the few-shot learning scenario.
+Weight imprinting is implemented by setting the final layer weights for the novel classes to the scaled average of the embedding vectors of their training samples and was first introduced by Qi et al.
+(2018) for the few-shot learning scenario.
 The authors find that for up to $20$ samples, using a combination of imprinting and fine-tuning outperforms other state-of-the-art methods, including nearest neighbor algorithms.
 In contract, in our work we do not limit the number of samples and perform no fine-tuning on the imprinted weights to maintain efficiency.
-Imprinting has been applied to object detection (imprint-objdetec; yan2023few), multi-label classification (imprint-multilabel), semantic segmentation (masked-imprinting), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (fewshot-metagenweights).
-itsdone apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights.
+Imprinting has been applied to object detection (Li et al., 2021; Yan et al., 2023), multi-label classification (Khan et al., 2021), semantic segmentation (Siam et al., 2019), and combined with an attention mechanism to generate weights for the novel classes in a few-shot classification task (Gidaris and Komodakis, 2018).
+Hosoda et al.
+(2024) apply imprinting using quantile normalization to ensure statistical similarity between new and existing weights.
 We consider this as one normalization scheme in our framework.
-covidimprint apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available.
+Zhang et al.
+(2021) apply imprinting in chest radiography for detection of COVID-19 and find that it yields better results than joint gradient descent training of all classes when only few samples are available.
 They speculate whether normalization is a constraint in their imprinting model.
-Before the era of deep learning, first-nearestclassmean analyze the transferability of hand-crafted image features.
+Before the era of deep learning, Mensink et al.
+(2013) analyze the transferability of hand-crafted image features.
 They use a “nearest class multiple centroids” (NCMC) classifier with multiple proxies generated from a $k$-means clustering algorithm.
@@
 Transfer Learning.
-Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (donahue_decaf_2014) and natural language processing (devlin_bert_2019).
-kornblith_better_2019 show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks.
-In addition, huh_what_2016 provide insights into the required quality of pre-training data.
+Using embedding vectors from pre-trained models is a simple and widely used transfer learning approach, established in the seminal works on computer vision (Donahue et al., 2014) and natural language processing (Devlin et al., 2019).
+Kornblith et al.
+(2019) show that pre-training performance of a model is highly correlated with the performance of the resulting embedding vectors in downstream tasks.
+In addition, Huh et al.
+(2016) provide insights into the required quality of pre-training data.
 Our work is orthogonal to these studies, since we focus on studying weight generation, normalization, and aggregation techniques applied later on for new task adaptation.
@@
 Although we investigate transfer learning scenarios, we review the imprinting applications and results from CL.
-icarl dynamically select a subset of examples for each class and update internal representations via gradient descent.
+Rebuffi et al.
+(2017) dynamically select a subset of examples for each class and update internal representations via gradient descent.
 They use a nearest mean classifier (NMC) with respect to the saved examples.
-ncmbaseline use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings.
+Janson et al.
+(2022) use an NMC classifier as well and achieve good performance on CL benchmarks without any fine-tuning of the embeddings.
 However, they do not investigate the effect of normalization and using multiple proxies.
-Findings of imprintallknn show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored.
+Findings of Prabhu et al.
+(2023) show that a simple, approximate $m$-nearest neighbor classifier outperforms existing methods in an Online CL setting when all data can be stored.
 In our work, however, we compare imprinting all data to a limited number of more representative proxies striving for efficiency.
 Neural Collapse (NC).
-The phenomenon of NC was identified by papyan2020prevalence and refers to the convergence of the last-layer weight vectors to class means.
-It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (zhu2021geometric; neuralcollapse2; neuralcollapse3).
-Nevertheless, complete neural collapse is practically unrealistic (DBLP:conf/icml/TirerHN23).
-In transfer learning, DBLP:conf/iclr/Galanti0H22 show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios.
+The phenomenon of NC was identified by Papyan et al.
+(2020) and refers to the convergence of the last-layer weight vectors to class means.
+It was shown that, regardless of the loss function, optimizer, batch-normalization, or regularization, NC will eventually occur (provided the training data has a balanced distribution) (Zhu et al., 2021; Han et al., 2022; Kothapalli, 2023).
+Nevertheless, complete neural collapse is practically unrealistic (Tirer et al., 2023).
+In transfer learning, Galanti et al.
+(2022) show that NC occurs on new samples and classes from the same distribution as the pre-training dataset, highlighting the usability of foundational models in such scenarios.
 In our work, we expand the survey on NC by experimenting with out-of-distribution classes belonging to different datasets and linking their degree of collapse to the success of certain imprinting strategies.
@@
 Hereby, embeddings from the FM are normalized before the generation (GEN) step according to NORMpre.
-The generated weight vectors per class are referred to as proxies, prototypes, or representatives (metric-proxies; prototypical; convolutional-prototypes; semanticdriftCL).
+The generated weight vectors per class are referred to as proxies, prototypes, or representatives (Movshovitz-Attias et al., 2017; Snell et al., 2017; Yang et al., 2018; Yu et al., 2020).
 These proxies are normalized according to NORMpost.
-As in the work by qi-imprinting, we do not use bias values.
+As in the work by Qi et al.
+(2018), we do not use bias values.
 To classify the test data in T during inference, it is first embedded by the FM, normalized according to NORMinf, and finally aggregated by AGG, resulting in a predicted class label.
@@
 The purpose of GEN is to determine how the embeddings of the training data in T are used to form the new weights.
-In contrast to qi-imprinting which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in first-nearestclassmean and enable non-linear classification.
+In contrast to Qi et al.
+(2018) which only incorporates one proxy per class (the mean), we add flexibility by allowing each class to have multiple proxies as in Mensink et al.
+(2013) and enable non-linear classification.
 We denote the number of proxies per class as $k$, ranging between $1$ and the number of samples, and investigate the following operations conducted per class to generate those:
@@
   • mean: The mean of all embeddings.
-  • k-means: $k$-means cluster centers using KMeans from sklearn (scikit-learn). $k=1$ is the same as mean.
+  • k-means: $k$-means cluster centers using KMeans from sklearn (Pedregosa et al., 2011). $k=1$ is the same as mean.
   • k-medoids: $k$-medoids cluster centers using KMedoids from sklearn.
@@
 In any case, the vector is $L^{2}$-normalized by dividing it by its Euclidean length $\|\cdot\|_{2}$.
-quantile normalization (quantnorm1; quantnorm2) can only be applied to generated weights.
+quantile normalization (Amaratunga and Cabrera, 2001; Bolstad et al., 2003) can only be applied to generated weights.
 This non-linear operation distributes weights equally.
@@
 In particular, for the first class there is no reference distribution to map to.
-This is different from itsdone, where new weights are matched to the distribution of the original classifier weights of the FM.
+This is different from Hosoda et al.
+(2024), where new weights are matched to the distribution of the original classifier weights of the FM.
 Since we do not consider the classes used for pre-training the FM and especially do not assume access to their last-layer weights, this is not possible in our scenario.
@@
 3.2 Quantifying Neural Collapse
-Neural collapse (NC) (papyan2020prevalence) characterizes the state of the features produced by a classification neural network after training to near zero training loss.
+Neural collapse (NC) (Papyan et al., 2020) characterizes the state of the features produced by a classification neural network after training to near zero training loss.
 Namely, the learned embeddings of each class converge, i.e., collapse, to their class means.
@@
 Two important characteristics of NC are variability collapse, i.e., the within-class variability of the penultimate-layer embeddings collapses to zero, and convergence to nearest-mean-classification.
-We focus on variability collapse ($\mathcal{NC}_{1}$) as in zhu2021geometric.
+We focus on variability collapse ($\mathcal{NC}_{1}$) as in Zhu et al.
+(2021).
 Given a foundation model FM and a finite dataset containing $C$ classes with (for simplicity) $N$ samples per class, we use its $l$-dimensional embeddings $\{\boldsymbol{\mathbf{h}}_{c,i}\}_{1\leq c\leq C,\;1\leq i\leq N}$ given by FM to define the global feature mean $\boldsymbol{\mathbf{h}}_{G}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, class means $\overline{\boldsymbol{\mathbf{h}}}_{c}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{\mathbf{h}}_{c,i}$, the within-class covariance matrix $\mathbf{\Sigma}_{W}\coloneqq\frac{1}{CN}\sum_{c=1}^{C}\sum_{i=1}^{N}\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)\bigl(\boldsymbol{\mathbf{h}}_{c,i}-\overline{\boldsymbol{\mathbf{h}}}_{c}\bigr)^{\top}$, and the between-class covariance matrix $\mathbf{\Sigma}_{B}\coloneqq\frac{1}{C}\sum_{c=1}^{C}\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)\bigl(\overline{\boldsymbol{\mathbf{h}}}_{c}-\boldsymbol{\mathbf{h}}_{G}\bigr)^{\top}$ to finally compute
@@
 Note that this measurement is not independent of the embedding dimension $l$ and the number of classes $C$.
-According to NC, imprinting the mean, as originally done by qi-imprinting, is best when $\mathcal{NC}_{1}$ is small.
+According to NC, imprinting the mean, as originally done by Qi et al.
+(2018), is best when $\mathcal{NC}_{1}$ is small.
 We claim that when the data is not fully collapsed (as is often the case in practice), the scale of $\mathcal{NC}_{1}$ could guide the proxy generation method, e.g., having multiple proxies $k>1$ per class.
@@
 Non-parametric rank-based tests are well suited for this setting.
-Following cddiag, we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric.
-Empirical evidence in cddiag further suggests that these tests have good power in classifier comparison scenarios.
+Following Demšar (2006), we use the Friedman test and Wilcoxon signed-rank tests because they do not assume normality or homoscedasticity, and can be applied uniformly to any evaluation metric.
+Empirical evidence in Demšar (2006) further suggests that these tests have good power in classifier comparison scenarios.
 Concretely, for each evaluation instance (i.e., foundation model FM and task T) we compute the classification accuracy of all configurations and assign ranks (rank 1 = highest accuracy).
@@
 Specifically, for every pair of configurations we run a two-sided Wilcoxon signed-rank test across evaluation instances.
-The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (holm1979simple).
+The resulting $p$-values are corrected for multiple comparisons using the Holm step-down procedure (Holm, 1979).
 Critical difference (CD) diagrams summarize both the average performance and the significance structure.
@@
 Foundation Models FM.
-We use resnet18, resnet50 (resnet), vit_b_16 (vit), and swin_b (swin) as FMs, two CNN-based and two Transformer-based architectures.
-All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (imagenet).
+We use resnet18, resnet50 (He et al., 2016), vit_b_16 (Dosovitskiy et al., 2021), and swin_b (Liu et al., 2021) as FMs, two CNN-based and two Transformer-based architectures.
+All four models are pre-trained on ImageNet-1K (ILSVRC 2012) (Deng et al., 2009).
 To generate the embeddings, we use PyTorch’s torchvision models.
@@
 For that, we randomly pre-sample the training data of T to $n$ samples per class – transitioning into the low-data regime.
-To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (mnist), FashionMNIST (fashionmnist), and CIFAR-10 (cifar10), each containing 10 classes.
+To find out the best imprinting strategy within our IMPRINT framework, we focus on tasks T created from the datasets MNIST (Deng, 2012), FashionMNIST (Xiao et al., 2017), and CIFAR-10 (Krizhevsky et al., 2009), each containing 10 classes.
 We mainly focus on the three T containing all ten classes.
@@
 See fig.˜4 for a simplified illustration.
-Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (mnistm), SVHN (svhn), and USPS (usps).
+Moreover, we construct a new out-of-distribution, non-collapsed dataset by merging all classes from four digit datasets: MNIST, MNIST-M (Ganin et al., 2016), SVHN (Netzer et al., 2011), and USPS (Hull, 1994).
 The resulting dataset, which we refer to as CombiDigits, exhibits reduced collapse due to the greater distributional diversity within each class.
@@
 We report the average rank and assess statistical significance of ranking (dis-)agreements using critical difference (CD) diagrams as explained in section˜3.3.
-The code used to generate these diagrams is inspired by cddiag-code.
+The code used to generate these diagrams is inspired by Ismail Fawaz et al.
+(2019).
 In experiments with neural collapse (section˜5.3), we investigate the four FMs on 100 random ImageNet tasks with remapped labels (as explained above) and the four tasks containing all of MNIST, FashionMNIST, CIFAR-10, and CombiDigits, respectively.
@@
 5.1 Best Imprinting Strategy
-We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by qi-imprinting; itsdone; ncmbaseline, as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT.
+We provide a comparison between existing memory-constrained methods used for imprinting on foundation models in fig.˜2, namely, the work by Qi et al.
+(2018); Hosoda et al.
+(2024); Janson et al.
+(2022), as well as a novel, best-performing configuration (“Ours”) that results from IMPRINT.
 Focusing on $k=20$, we find that our method, consisting of k-means weight generation, L2 normalizations, and max aggregation, outperforms all previously studied approaches by a margin of $4\%$ on average with statistical significance.
@@
 Learned Weights.
-Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by harun2025good, which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here.
+Employing gradient-based methods – such as setting weights via the analytical least-squares initialization proposed by Harun and Kanan (2025), which uses data and labels from all classes jointly – can be seen as an upper bound (“Oracle” in fig.˜2) for the unsupervised weight imprinting approaches discussed here.
 We extend this method by combining it with k-means, forming k-least-squares, and observe in section˜A.5 that using multiple proxies per class instead of a single weight vector can still improve performance, particularly on datasets with high $\mathcal{NC}_{1}$.
@@
 However, the performance of NORMpre and NORMinf remains statistically indifferent for L2 weight normalization (see fig.˜A.2 for all combinations at once with $k=1$, and fig.˜A.3 for $k=20$).
-We restrict all the subsequent experiments to using L2 across all NORM following qi-imprinting.
+We restrict all the subsequent experiments to using L2 across all NORM following Qi et al.
+(2018).
 This combination of normalizations is chosen to specifically model cosine similarity within max aggregation.
@@
 In particular, we find that the mean weight generation (GEN) method, despite its prominence in previous work, falls short compared to k-means – even when the number of proxies $k$ is very small.
-Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by qi-imprinting, highlighting its advantage in low-data regimes.
+Remarkably, with as little as 50 samples per class, k-means can already outperform the original imprinting method proposed by Qi et al.
+(2018), highlighting its advantage in low-data regimes.
 L2 weight normalization is essential for strong performance.
@@
 Future Work.
-The usage of both weight and activation sparsity as shen2023reducing could change the within- and between-class variability in favor of using a higher number of proxies.
+The usage of both weight and activation sparsity as Shen et al.
+(2023) could change the within- and between-class variability in favor of using a higher number of proxies.
 Besides only using the penultimate layer embeddings for generating the classifier weights, an interesting area of study could be extracting embeddings from previous layers of the FM for this purpose.
-Recently, the study by marczak2025revisiting showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable.
-Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (masked-imprinting).
+Recently, the study by Marczak et al.
+(2025) showed that adding a multi-layer perceptron projector between the penultimate and classification layers results in representations that are more transferable.
+Another avenue of research is to thoroughly analyze imprinting the weights of other layers as well (Siam et al., 2019).
 Acknowledgments
@@
 We briefly describe the datasets used in our experiments.
-ImageNet (imagenet).
+ImageNet (Deng et al., 2009).
 We use the ILSVRC 2012 version (commonly called ImageNet-1K) containing 1 000 classes and 1.2M training images.
@@
 For neural collapse investigations, we construct tasks by randomly grouping $d$ classes into one label, producing “$d$ in 1” tasks as explained in section˜4.
-MNIST (mnist).
+MNIST (Deng, 2012).
 A benchmark dataset of handwritten digits (0–9), consisting of 60 000 training and 10 000 test grayscale images of size $28\times 28$.
-FashionMNIST (fashionmnist).
+FashionMNIST (Xiao et al., 2017).
 A drop-in replacement for MNIST with the same format and number of samples, containing grayscale images of fashion items across 10 classes.
-CIFAR-10 (cifar10).
+CIFAR-10 (Krizhevsky et al., 2009).
 A dataset of $32\times 32$ RGB images covering 10 object classes, with 50 000 training and 10 000 test samples.
-MNIST-M (mnistm), SVHN (svhn), USPS (usps).
+MNIST-M (Ganin et al., 2016), SVHN (Netzer et al., 2011), USPS (Hull, 1994).
 Digit classification datasets each containing digits 0-9 with domain-specific visual characteristics.
@@
 A.3 Imprinting as Memory
-We revisit the classical idea of Bidirectional Associative Memories (BAMs) (kohonen; anderson72; nakano; anderson77) and the associated update rule (sejnowski; dayan) for storing key-value pairs as in
+We revisit the classical idea of Bidirectional Associative Memories (BAMs) (Kohonen, 2009; Anderson, 1972; Nakano, 2007; Anderson et al., 1977) and the associated update rule (Sejnowski, 1977; Dayan and Willshaw, 1991) for storing key-value pairs as in
 $\mathbf{W}\leftarrow\mathbf{W}+\boldsymbol{\mathbf{v}}\boldsymbol{\mathbf{k}}^{\top},$
@@
 Notably, this form of direct memory update has seen renewed attention in modern architectures beyond standard query-key-value attention.
-In particular, the xLSTM model (xlstm) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
+In particular, the xLSTM model (Beck et al., 2024) implements this mechanism within its mLSTM memory blocks, where the matrix memory cell is updated by gated key-value associations, closely following this classical covariance rule, indicating a broader resurgence of associative memory principles in contemporary sequence modeling.
 A.4 Differences between Foundation Models
@@
 This seemingly minor change reveals a striking contrast between CNN- and Transformer-based architectures: a distinct and consistent dip between $k=1$ and $k=d$ appears in Transformer-based models, whereas this dip is absent in fig.˜11(b), where L2 is used as NORMpre, and does not occur at all in the resnet models.
-We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., itsdone).
+We hypothesize that this difference arises from the distinct embedding distributions of CNN- and Transformer-based architectures (see, e.g., Hosoda et al.
+(2024, Figure S2)).
 A.5 Learned Weights (with Multiple Proxies) and Comparisons
-harun2025good recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions.
+Harun and Kanan (2025) recently studied the initialization of classifier weights for novel categories in a Continual Learning (CL) setting using the least squares algorithm, comparing it to random initialization and class mean imprinting across various loss functions.
 In contrast, our work centers on imprinting-based approaches, which avoid using gradient-based optimization and cross-class statistics, operating instead on a per-class basis with immediate availability of data.
@@
 Model $\lambda$
-resnet18, resnet50 (resnet) 0.0001
-vit_b_16 (vit) 0.1
-swin_b (swin) 0.05
+resnet18, resnet50 (He et al., 2016) 0.0001
+vit_b_16 (Dosovitskiy et al., 2021) 0.1
+swin_b (Liu et al., 2021) 0.05
 Equation˜A.1 reveals that least-squares shares structural similarities with mean imprinting, but applies a more sophisticated normalization scheme that includes both scaling and rotation.